### v1

# Conduct Parameter Estimation in Homogeneous Goods Markets with Equilibrium Existence and Uniqueness Conditions: The Case of Log-linear Specification

Abstract
We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in a log-linear model with homogeneous goods markets. First, we derive such conditions. Second, Monte Carlo simulations confirm that in a log-linear model, incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters.
Keywords: Conduct parameters, Homogenous Goods Market, Mathematical Programming with Equilibrium Constraints, Monte Carlo simulation
JEL Codes: C5, C13, L1
1 Introduction
Measuring competitiveness is a crucial task in the empirical industrial organization literature. Conduct parameter is considered a useful measure of competitiveness. However, it cannot be directly measured from data because data usually lack information about marginal cost. Therefore, researchers aim to identify and estimate the conduct parameter.
As the simplest specification, Bresnahan (1982) considers identification of conduct parameters for the linear model. Matsumura and Otani (2023b, a) resolves the conflict on some identification problems between Bresnahan (1982) and Perloff and Shen (2012). However, researchers often implement alternative specifications, such as the log-linear model (Okazaki et al. 2022, Mérel 2009). In the context of the log-linear model, these papers identify estimation issues, with some estimated conduct parameters being unrealistically low or even negative. This raises doubts about the methodology, despite the identification strategy outlined by Lau (1982). This is an obstacle to choosing a better specification of the demand and supply functions.
To overcome the problem, we propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices as constraints. First, we prove that a unique equilibrium exists under certain theoretical conditions. As far as we know, this is a new result. Second, we show that incorporating equilibrium existence and uniqueness conditions resolves the above estimation problems and makes the model work in estimating conduct parameters even in the log-linear model. Hence, our results support those of Lau (1982) numerically.
2 Model
Consider data with $T$ markets with homogeneous products. Assume there are $N$ firms in each market. Let $t=1,\ldots,T$ be the index of markets. Then, we obtain the supply equation:
$P_{t}=-\theta\frac{\partial P_{t}(Q_{t})}{\partial Q_{t}}Q_{t}+MC_{t}(Q_{t}),$ (1)
where $Q_{t}$ is the aggregate quantity, $P_{t}(Q_{t})$ is the demand function, $MC_{t}(Q_{t})$ is the marginal cost function, and $\theta\in[0,1]$, which is the conduct parameter. The equation nests perfect competition $(\theta=0)$, Cournot competition $(\theta=1/N)$, and perfect collusion $(\theta=$ 1)1)1 ) (Bresnahan 1982).
Consider an econometric model. Assume that the demand and the marginal cost functions are written as follows:
$P_{t}=f(Q_{t},X^{d}_{t},\varepsilon^{d}_{t},\alpha),$
$MC_{t}=g(Q_{t},X^{c}_{t},\varepsilon^{c}_{t},\gamma),$
where $X^{d}_{t}$ and $X^{c}_{t}$ are the vector of exogenous variables, $\varepsilon^{d}_{t}$ and $\varepsilon^{c}_{t}$ are the error terms, and $\alpha$ and $\gamma$ are the vector of parameters. We also have the demand- and supply-side instruments, $Z^{d}_{t}$ and $Z^{c}_{t}$, and assume that the error terms satisfy the mean independence condition $E[\varepsilon^{d}_{t}\mid X^{d}_{t},Z^{d}_{t}]=E[\varepsilon^{c}_{t}\mid X^{c}_{t},Z^{c}_{t}]=0$.
2.1 Log-linear demand and log-linear marginal cost
Consider a log-linear model, which is a typical specification. The demand and marginal cost functions are specified as
$\log P_{t}$ $=\alpha_{0}-(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t},$ (2)
$\log MC_{t}$ $=\gamma_{0}+\gamma_{1}\log Q_{t}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t},$ (3)
where $Y_{t}$ is an excluded demand shifter, $W_{t}$ and $R_{t}$ are excluded cost shifters, and $Z_{t}^{R}$ is Bresnahan’s demand rotation instrument. Then, Equation (1) is written as
$P_{t}$ $=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+MC_{t}.$ (4)
By taking logarithm of Equation (4) and substituting Equation (3), we obtain
$\log P_{t}=-\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))+\gamma_{0}+\gamma_{1}\log Q_{t}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}.$ (5)
In this model, the number of equilibrium prices varies. Although it is a widely known model, to our knowledge, there is no paper examining this. Thus, we derive the conditions for the uniqueness in the following proposition which are used in our estimation.
Proposition 1.
Assume that $\alpha_{1}+\alpha_{2}Z^{R}_{t}\neq 0$. Let $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}$. The number of equilibrium prices $P_{t}^{*}>0$ is determined as follows:
  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, there is no equilibrium price,
  • When $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0,
    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\neq 1$, there is a unique equilibrium price,
    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, there are infinitely many equilibrium prices when $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, but there is no equilibrium price otherwise.
See the online appendix A.1 for the proof.
3 Estimation
To estimate parameters in the demand and supply equations, we use GMM estimation. Among GMM estimators, we apply the nonlinear system two-stage-least-squares (N2SLS) using Equation (2) and (5).
Let $\xi=(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\theta)$ be the vector of parameters. Given the demand equation (2) and the supply equation (5), we can write the error terms in the demand and supply equation as
${\varepsilon}_{t}^{d}(\xi)$ $=\log P_{t}-\alpha_{0}+(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}-\alpha_{3}\log Y_{t},$ (6)
${\varepsilon}_{t}^{c}(\xi)$ $=\log P_{t}+\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))-\gamma_{0}-\gamma_{1}\log Q_{t}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t}.$ (7)
To estimate the parameters, we convert the conditional moment conditions, $E[\varepsilon_{t}^{d}\mid Z_{t}^{d}]=E[\varepsilon_{t}^{c}\mid Z_{t}^{c}]=0$, into unconditional moment conditions, $E[\varepsilon_{t}^{d}Z_{t}^{d}]=E[\varepsilon_{t}^{c}Z_{t}^{c}]=0$. Using Equations (6) and (7), we construct the sample analog of the moment conditions:
$g(\xi)=\left[\begin{array}[]{l}\frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{d}_{t}(\xi)Z_{t}^{d}\\ \frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{c}_{t}(\xi)Z_{t}^{c}\end{array}\right].$
We define the GMM estimator as the vector $\xi^{*}$ that solves the problem,
$\min_{\xi}\ g(\xi)^{\top}Wg(\xi)$ (8)
where the weight matrix $W$ is defined as
$W=\left[\frac{1}{T}\sum_{t=1}^{T}Z_{t}^{\top}Z_{t}\right]^{-1}\text{ where }Z_{t}=\left[\begin{array}[]{ll}Z_{t}^{d\top}&0\\ 0&Z_{t}^{c\top}\end{array}\right].$ (11)
The solution to (8) is called the N2SLS estimator (Wooldridge 2010).
Using Proposition 1, we impose the following constraints:
$0\leq\theta\leq 1,$ (12)
$\alpha_{1}+\alpha_{2}Z_{t}^{R}>0,\quad\gamma_{1}>0,\quad t=1,\ldots,T$ (13)
$1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})>0,\quad t=1,\ldots,T.$ (14)
Constraint (12) is a standard assumption on the conduct parameter. Constraints (13) and (14) relate to the uniqueness of equilibrium prices. Constraint (13) implies the downward-sloping demand and upward-sloping marginal cost. See online appendix A.2 for the detailed setting.
4 Simulation results
We compare N2SLS estimation with and without Constraints (12), (13), and (14). Table 1 presents the results. First, as reported in the literature (Mérel 2009, Okazaki et al. 2022), N2SLS without constraints converges to implausibly low and negative conduct parameters. This is because the algorithm searches the area in which no equilibrium price and quantity exist and then finds unreasonable local optima that might have a lower objective function than at the true values. Second, N2SLS with constraints converges to the reasonable values. As the sample size increases, the bias and root-mean-square error (RMSE) decrease to levels comparable to linear models. Third, we find an ad hoc method that uses Equation (3) to compute the residual in the supply estimation, $\varepsilon_{t}^{c}$, and Equation (4) as an equality constraint with Constraints (12), (13), and (14). Then, the bias and RMSE of conduct parameter $\theta$ are reduced to 0.014 and 0.217 in Panel (c), although the results do not dominate the results in Panel (b) for all parameters. Therefore, incorporating equilibrium existence and uniqueness conditions simply is helpful for conduct parameter estimation in nonlinear models such as a log-linear model. See online appendix A.4 for additional experiments under different variances of errors and the results of a linear model.
Table 1: Performance comparison
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$1.070 7.012 -0.021 5.110 0.365 2.207 0.400 2.030
$\alpha_{1}$0.164 1.060 -0.001 0.782 0.073 0.458 0.096 0.574
$\alpha_{2}$0.011 0.104 -0.006 0.071 0.002 0.033 0.005 0.043
$\alpha_{3}$0.101 0.619 -0.005 0.474 0.021 0.198 0.029 0.187
$\gamma_{0}$ 9.735 15.743 9.636 10.870 13.173 13.269 13.294 13.351
$\gamma_{1}$0.070 1.624 -0.177 0.469 -0.184 0.248 -0.177 0.220
$\gamma_{2}$0.034 0.939 -0.098 0.317 -0.090 0.152 -0.080 0.127
$\gamma_{3}$0.047 0.750 -0.091 0.311 -0.098 0.156 -0.085 0.133
$\theta$3e+05 3e+06 -2e+05 2e+06 -8e+04 9e+04 -9e+04 1e+05
Runs converged (%) 99.500 99.800 98.600 98.400
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.905 6.954 0.120 5.001 0.072 2.042 0.052 1.563
$\alpha_{1}$0.141 1.053 0.018 0.768 0.010 0.313 0.008 0.241
$\alpha_{2}$0.006 0.101 0.000 0.062 0.001 0.024 0.001 0.019
$\alpha_{3}$0.088 0.620 0.007 0.475 -0.001 0.186 0.003 0.152
$\gamma_{0}$1.748 14.206 -0.938 6.428 0.015 1.995 0.163 1.570
$\gamma_{1}$ 0.254 1.927 0.129 0.825 0.018 0.226 0.003 0.170
$\gamma_{2}$ 0.117 1.083 0.049 0.467 0.008 0.148 0.007 0.116
$\gamma_{3}$ 0.098 0.890 0.058 0.478 -0.001 0.148 0.003 0.118
$\theta$0.100 0.441 -0.072 0.424 -0.121 0.351 -0.148 0.333
Runs converged (%) 99.600 99.900 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Note: The error terms are drawn from a normal distribution, $N(0,\sigma)$. True values: $\alpha_{0}=20.0,\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=1.0,\gamma_{0}=5.0,\gamma_{1}=1.0,\gamma_{2}=1.0,\gamma_{3}=1.0,\theta=0.5$ and $\sigma=1.0$. See online appendix A.2 for the setting.
5 Conclusion
We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in homogeneous goods markets. First, we derive the conditions. Second, Monte Carlo simulations confirm that incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters in a log-linear model, as reported in the literature.
Acknowledgments
We thank Jeremy Fox, Yelda Gungor, and Isabelle Perrigne for valuable comments. This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Appendix A Online appendix
A.1 Existence and uniqueness of equilibrium prices
Proposition 1 proposes the conditions for the unique existence of $P_{t}(>0)$ solving the demand equation (2) and supply equation (5) for $P_{t}$ under $\theta\in[0,1]$. The proof is not based on the optimization by individual firms but checks if there exists a point at which the demand function and the marginal cost function cross. Therefore, we allow an equilibrium price to exist when the demand and the marginal cost are upward-sloping.
Proof.
Rewriting the demand equation (2) as
$\log Q_{t}(P_{t})=\frac{\alpha_{0}-\log P_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{(\alpha_{1}+\alpha_{2}Z^{R}_{t})}$
and substituting this into the supply equation (5), we obtain
$P_{t}$ $=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp\left(\gamma_{0}+\gamma_{1}\log Q_{t}(P_{t})+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}\right).$
$=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp\left(\gamma_{0}+\gamma_{1}\frac{\alpha_{0}-\log P_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{(\alpha_{1}+\alpha_{2}Z^{R}_{t})}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}\right)$
$=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp\left(\Xi+\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\log P_{t}\right)$
$=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ (15)
where $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}$.
Any price that satisfies (15) becomes an equilibrium price. To find an equilibrium price $P_{t}^{*}$, we define $\Delta(P_{t})$ as follows:
$\Delta(P_{t})$ $=\underbrace{[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}}_{(I)}-\underbrace{\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}}_{(II)}.$ (16)
We label the first term as (I) and the second term as (II).
An important note is that $P_{t}=0$ satisfies $\Delta(P_{t})=0$. Thus $P_{t}^{*}=0$ holds. However, our interest is a unique positive equilibrium price, so we seek conditions in which there is a unique positive price that satisfies $\Delta(P_{t})>0$.
When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) is always negative on $P_{t}>0$. In contrast, (II) is non-negative regardless of the sign of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ on $P_{t}>0$. Therefore, $\Delta(P_{t})$ is always negative on $P_{t}>0$, which implies that there is no positive equilibrium price.
When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) becomes a line passing through the origin with a positive slope, illustrated in Figure 1. Note that the first and second derivatives of (II) are given as
$\frac{d}{dP_{t}}\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ $=-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\exp(\Xi)P_{t}^{-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1},$ (17)
$\frac{d^{2}}{dP_{t}^{2}}\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ $=-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\left(-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1\right)\exp(\Xi)P_{t}^{-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-2}.$ (18)
Therefore, the shape of (II) on $P_{t}>0$ is changed based on the value of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ and the number that (I) and (II) cross on $P_{t}>0$ is also determined.
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})<0$ is illustrated in Figure 1(a). In this case, (II) is a monotone decreasing convex function because the first and second derivatives are both negative. As (I) is monotone increasing in $P_{t}>0$, (I) and (II) cross only once on $P_{t}>0$.
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\in[0,1)$ is illustrated in Figure 1(b). In this case, (II) becomes a monotone increasing concave function passing through the origin because the first derivative is positive and the second derivative is negative. As the value of the first derivative is infinity at $P_{t}=0$ since $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})-1<0$, (II) should be greater than (I) around $P_{t}=0$. Then as $P_{t}$ becomes large starting from zero, the difference in both terms becomes small, and both terms cross only once on $P_{t}>0$ eventually.
When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, (16) becomes
$0=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)]P_{t}.$
If $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})=\exp(\Xi)$, the above equation holds for any $P_{t}$. Thus there are infinitely many equilibrium prices on $P_{t}>0$, which is illustrated in Figure 1(c). When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq\exp(\Xi)$, there is no equilibrium price on $P_{t}>0$ because the right-hand side of the equation is always non-zero on $P_{t}>0$, which is illustrated in Figure 1(d).
The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})>1$ is illustrated in Figure 1(e). In this case, (II) becomes a monotone increasing convex function passing through the origin because the first and second derivatives are positive. Therefore, (I) and (II) cross only once on $P_{t}>0$.
Figure 1: The illustrations of the equilibrium existence and uniqueness conditions
$P_{t}$$\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$$[1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})]P_{t}$
(a) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}<0$
$P_{t}$
(b) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\in[0,1)$
$P_{t}$
(c) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}=1,1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})=\exp(\Xi)$
$P_{t}$
(d) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}=1,1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})\neq\exp(\Xi)$
$P_{t}$
(e) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}>1$
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. The thick line represents the first term and the dotted line represents the second term in (16).
From $\Delta(P_{t})=0$, the positive equilibrium price can be written as
00 0 $=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$
00 0 $=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$
$P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$ $=\frac{1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})}{\exp(\Xi)}$
$P_{t}^{*}$ $=\left(\frac{1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})}{\exp(\Xi)}\right)^{-\frac{\alpha_{1}+\alpha_{2}Z^{R}_{t}}{\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}_{t}}}.$ (19)
Figure 2 illustrates how the demand and supply equations cross under the conditions. When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the supply equation (5) is ill-defined because the inside of the log function becomes negative. Thus, there should not be any equilibrium. Hereafter, assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand equation and supply equation coincides. Thus there are infinitely many equilibria in the model. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand and supply equations have a same slope and different intercepts, which means that both equations become parallel. Thus, there is no equilibrium. When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, the demand and supply equations have different slopes, and hence we can find a unique equilibrium.
Figure 2: The illustrations of the equilibrium existence and uniqueness conditions
[…]
(a) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
[…]
(b) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
[…]
(c) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$, and $\gamma_{1}>0$
[…]
(d) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}\leq 0$, and $\gamma_{1}>0$
[…]
(e) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$, and $\gamma_{1}\leq 0$
[…]
(f) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}\leq 0$, and $\gamma_{1}\leq 0$
Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0. Case (a) and (b) also allow both the demand and supply equations to be downward sloping.
A.2 Simulation and estimation procedure
To generate the simulation data, for each model, we first generate the exogenous variables $Y_{t},Z^{R}_{t},W_{t},R_{t},H_{t}$, and $K_{t}$ and the error terms $\varepsilon_{t}^{c}$ and $\varepsilon_{t}^{d}$ based on the data generation process in Table 2. By substituting the Equation (2) into Equation (5) and solving it for $P_{t}$, the log aggregate quantity is given as:
$\log Q_{t}$ $=\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))-\gamma_{0}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t}+\varepsilon^{d}_{t}-\varepsilon^{c}_{t}}{\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}_{t}}.$ (20)
We compute the equilibrium quantity $Q_{t}$ for the log-linear model by (20). We then compute the equilibrium price $P_{t}$ by substituting $Q_{t}$ and other variables into the demand function (2). We generate 1000 data sets of 100, 200, 1000, 1500 markets. We jointly estimate the demand and supply parameters by the simultaneous equation model (Wooldridge 2010) from the true values. We use state-of-the-art constrained optimization solvers, i.e., Ipopt.jl which implements an interior point line search filter method that aims to find a local solution of nonlinear programming problems.
Table 2: True parameters and distributions
$\alpha_{0}$ $20.0$
$\alpha_{1}$ $1.0$
$\alpha_{2}$ $0.1$
$\alpha_{3}$ $1.0$
$\gamma_{0}$ $5.0$
$\gamma_{1}$ $1.0$
$\gamma_{2}$ $1.0$
$\gamma_{3}$ $1.0$
$\theta$ $0.5$
Demand shifter
$Y_{t}$ $N(0,1)$
Demand rotation instrument
$Z^{R}_{t}$ $U(0,1)$
Cost shifter
$W_{t}$ $U(1,3)$
$R_{t}$ $U(1,3)$
$H_{t}$ $W_{t}+U(0,1)$
$K_{t}$ $R_{t}+U(0,1)$
Error
$\varepsilon^{d}_{t}$ $N(0,\sigma)$
$\varepsilon^{c}_{t}$ $N(0,\sigma)$
Note: $\sigma=\{0.5,1.0,2.0\}$. $N:$ Normal distribution. $U:$ Uniform distribution.
A.3 Why does the estimation without equilibrium conditions converge to the extremely low conduct parameters?
We should interpret reasons why the estimation without equilibrium conditions converges to the negative conduct parameters, as shown in Panel (a) in Table 1, although the obtained parameters cannot be interpreted as the exogenous elements determining equilibrium outcomes. First, we rewrite the terms in Equation (7) into
$\log(1-\theta C_{t})-\gamma_{0}$
where $C_{t}=\alpha_{1}+\alpha_{2}Z^{R}_{t}$ is given in the demand estimation. If the model allows $\theta<0$, then $1-\theta C_{t}$ can take an arbitrarily large value through log-transformation relative to $\gamma_{0}$. For example, if $0001-\theta C_{t}=1,000,$ 000, then $\log(1-\theta C_{t})=13.8$, which has a small contribution relative to $\gamma_{0}$ for evaluating the moment conditions.
A.4 Additional experiments
Additional results for different $\sigma$ are shown in Tables 3 and 4. As a summary, the main findings in the main text are robust.
In Tables 6 and 7, we illustrate that N2SLS with Constraints (12) works for the linear model as in Matsumura and Otani (2023b). This means that incorporating the conditions is innocuous for the linear model.
Additional ad hoc improvement is possible by using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14). Table 5 shows that the estimation of the conduct parameter $\theta$ improves, that is, the bias is 0.014 and RMSE is 0.217. However, the results with ad hoc improvement do not dominate the results without the improvement for all parameters.
Table 3: Performance comparison ($\sigma=0.5$)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$ 0.119 4.055 0.346 3.180 0.344 1.497 0.379 1.582
$\alpha_{1}$ 0.019 0.620 0.054 0.486 0.076 0.429 0.099 0.576
$\alpha_{2}$ 0.000 0.050 0.003 0.039 0.002 0.029 0.003 0.036
$\alpha_{3}$ 0.009 0.336 0.025 0.268 0.028 0.129 0.031 0.129
$\gamma_{0}$ 10.104 11.066 9.397 10.288 13.240 13.280 13.314 13.346
$\gamma_{1}$0.146 0.340 -0.166 0.264 -0.180 0.198 -0.181 0.192
$\gamma_{2}$0.072 0.229 -0.079 0.165 -0.087 0.108 -0.086 0.099
$\gamma_{3}$0.069 0.230 -0.082 0.171 -0.088 0.110 -0.087 0.100
$\theta$5e+04 1e+05 -3e+04 6e+04 -8e+04 9e+04 -8e+04 9e+04
Runs converged (%) 100.000 100.000 98.100 98.700
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.006 3.556 0.175 3.158 0.059 0.977 -0.017 0.763
$\alpha_{1}$0.002 0.545 0.026 0.484 0.009 0.150 -0.003 0.117
$\alpha_{2}$ 0.002 0.044 0.004 0.037 0.000 0.012 0.000 0.010
$\alpha_{3}$ 0.001 0.309 0.012 0.266 0.005 0.091 0.000 0.073
$\gamma_{0}$0.616 3.317 -0.221 2.191 0.275 1.076 0.421 0.972
$\gamma_{1}$ 0.089 0.399 0.038 0.252 -0.015 0.107 -0.027 0.089
$\gamma_{2}$ 0.044 0.265 0.022 0.170 -0.007 0.074 -0.011 0.061
$\gamma_{3}$ 0.046 0.273 0.019 0.178 -0.008 0.078 -0.012 0.059
$\theta$0.064 0.390 -0.079 0.339 -0.119 0.260 -0.145 0.261
Runs converged (%) 99.900 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Table 4: Performance comparison ($\sigma=2.0$)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$2.358 11.286 -1.457 8.696 0.231 4.747 0.450 4.028
$\alpha_{1}$0.360 1.770 -0.227 1.313 0.040 0.744 0.095 0.711
$\alpha_{2}$0.021 0.138 -0.012 0.142 0.002 0.063 0.006 0.065
$\alpha_{3}$0.175 0.967 -0.138 0.867 0.006 0.434 0.031 0.347
$\gamma_{0}$ 9.868 21.289 10.490 13.359 13.151 13.453 13.177 13.384
$\gamma_{1}$0.015 2.542 -0.130 1.042 -0.202 0.400 -0.186 0.339
$\gamma_{2}$0.057 1.656 -0.083 0.735 -0.090 0.274 -0.099 0.224
$\gamma_{3}$0.020 2.115 -0.065 0.719 -0.097 0.279 -0.092 0.232
$\theta$7e+05 5e+06 -3e+05 2e+06 -1e+05 4e+05 -9e+04 1e+05
Runs converged (%) 99.600 99.800 98.300 98.900
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$2.008 11.292 -1.149 8.680 0.145 4.655 0.274 3.838
$\alpha_{1}$0.308 1.773 -0.181 1.313 0.020 0.713 0.041 0.590
$\alpha_{2}$0.016 0.136 -0.007 0.139 0.003 0.057 0.004 0.043
$\alpha_{3}$0.147 0.985 -0.113 0.875 -0.003 0.429 0.017 0.333
$\gamma_{0}$ 0.661 11.910 -1.090 13.475 -0.665 4.501 -0.404 3.245
$\gamma_{1}$0.044 1.600 0.172 1.770 0.098 0.556 0.066 0.391
$\gamma_{2}$0.056 1.086 0.088 1.172 0.054 0.346 0.024 0.252
$\gamma_{3}$0.020 1.205 0.108 1.225 0.049 0.365 0.031 0.267
$\theta$0.229 0.475 -0.163 0.464 -0.090 0.410 -0.108 0.394
Runs converged (%) 99.200 99.200 99.700 100.000
Sample size ($T$) 100 200 1000 1500
Table 5: Ad hoc method using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14)
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.347 2.978 -0.153 2.547 0.065 0.976 -0.009 0.760
$\alpha_{1}$0.043 0.452 -0.016 0.388 0.010 0.150 -0.001 0.117
$\alpha_{2}$0.024 0.065 -0.017 0.054 0.000 0.012 0.000 0.010
$\alpha_{3}$0.028 0.265 -0.013 0.224 0.005 0.091 0.001 0.073
$\gamma_{0}$2.458 5.660 -1.258 2.959 -0.132 0.942 -0.153 0.748
$\gamma_{1}$ 0.112 0.602 0.057 0.260 0.014 0.103 0.014 0.082
$\gamma_{2}$ 0.053 0.314 0.030 0.172 0.007 0.073 0.010 0.060
$\gamma_{3}$ 0.061 0.426 0.028 0.181 0.006 0.077 0.008 0.058
$\theta$ 0.169 0.407 0.109 0.356 -0.003 0.155 0.017 0.116
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.835 9.908 -0.561 6.310 0.147 4.424 0.301 3.686
$\alpha_{1}$0.109 1.563 -0.076 0.935 0.024 0.677 0.047 0.566
$\alpha_{2}$0.043 0.124 -0.033 0.135 -0.005 0.061 0.001 0.045
$\alpha_{3}$0.047 0.909 -0.078 0.664 -0.004 0.404 0.018 0.323
$\gamma_{0}$6.629 20.418 -5.373 16.692 -1.445 4.738 -1.123 3.411
$\gamma_{1}$ 0.422 2.630 0.384 2.103 0.137 0.580 0.115 0.409
$\gamma_{2}$ 0.184 1.788 0.185 1.318 0.073 0.357 0.048 0.261
$\gamma_{3}$ 0.199 1.885 0.215 1.496 0.068 0.376 0.055 0.276
$\theta$ 0.225 0.459 0.227 0.447 0.098 0.357 0.084 0.313
Runs converged (%) 99.200 99.800 99.800 100.000
Sample size ($T$) 100 200 1000 1500
Table 6: N2SLS without Constraints (12) Results of the linear model
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.018 0.465 0.007 0.323 -0.008 0.213 -0.006 0.097
$\alpha_{1}$0.045 2.257 0.024 1.523 0.018 1.016 -0.031 0.455
$\alpha_{2}$0.001 0.255 -0.001 0.176 -0.004 0.115 0.001 0.051
$\alpha_{3}$0.005 0.108 0.003 0.075 -0.001 0.050 -0.001 0.022
$\gamma_{0}$0.061 0.732 -0.005 0.474 -0.021 0.346 -0.005 0.152
$\gamma_{1}$0.311 3.450 -0.124 1.928 -0.081 1.303 -0.003 0.548
$\gamma_{2}$ 0.009 0.109 -0.001 0.071 0.003 0.051 0.000 0.023
$\gamma_{3}$ 0.001 0.108 0.003 0.075 0.003 0.053 0.000 0.022
$\theta$ 0.047 0.354 0.017 0.209 0.014 0.135 0.003 0.058
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.027 1.023 -0.002 0.641 -0.004 0.448 -0.016 0.188
$\alpha_{1}$0.024 4.396 -0.169 2.965 0.061 2.060 0.011 0.905
$\alpha_{2}$0.006 0.494 0.016 0.325 -0.007 0.234 -0.006 0.100
$\alpha_{3}$0.006 0.223 0.002 0.153 -0.001 0.099 -0.003 0.045
$\gamma_{0}$0.318 1.769 -0.091 1.059 -0.086 0.714 -0.004 0.308
$\gamma_{1}$ 5.859 210.853 -0.679 6.280 -0.338 2.972 -0.050 1.110
$\gamma_{2}$ 0.035 0.247 0.011 0.157 0.011 0.104 0.000 0.045
$\gamma_{3}$ 0.045 0.250 0.012 0.150 0.010 0.104 0.002 0.045
$\theta$0.399 18.450 0.098 0.738 0.054 0.308 0.009 0.113
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.263 2.596 0.071 1.670 -0.040 0.947 -0.002 0.412
$\alpha_{1}$0.271 10.820 0.008 6.492 0.236 4.263 0.021 1.809
$\alpha_{2}$0.044 1.253 0.023 0.779 -0.031 0.483 -0.003 0.210
$\alpha_{3}$0.024 0.584 0.008 0.343 -0.004 0.225 0.003 0.092
$\gamma_{0}$2.074 19.624 -0.551 3.043 -0.171 1.516 -0.051 0.633
$\gamma_{1}$ 58.209 1750.688 -2.416 56.909 -3.617 39.044 -0.103 2.334
$\gamma_{2}$ 0.242 2.430 0.065 0.409 0.020 0.220 0.006 0.093
$\gamma_{3}$ 0.230 2.328 0.055 0.404 0.010 0.219 0.008 0.092
$\theta$6.668 233.851 0.372 6.334 0.418 3.820 0.024 0.245
Sample size ($T$) 50 100 200 1000
Note: The data generating process follows Matsumura and Otani (2023b).
Table 7: N2SLS with Constraints (12) for the linear model
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.013 0.462 0.008 0.322 -0.008 0.213 -0.006 0.097
$\alpha_{1}$0.096 2.201 0.015 1.511 0.018 1.016 -0.031 0.455
$\alpha_{2}$ 0.006 0.247 0.001 0.174 -0.004 0.115 0.001 0.051
$\alpha_{3}$0.004 0.108 0.003 0.074 -0.001 0.050 -0.001 0.022
$\gamma_{0}$0.054 0.724 -0.002 0.472 -0.021 0.346 -0.005 0.152
$\gamma_{1}$0.098 2.620 -0.093 1.847 -0.081 1.303 -0.003 0.548
$\gamma_{2}$ 0.008 0.108 -0.002 0.070 0.003 0.051 0.000 0.023
$\gamma_{3}$ 0.001 0.107 0.003 0.075 0.003 0.053 0.000 0.022
$\theta$ 0.023 0.258 0.014 0.197 0.014 0.135 0.003 0.058
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$ 0.012 1.015 0.012 0.636 0.001 0.446 -0.016 0.188
$\alpha_{1}$0.383 4.216 -0.291 2.872 0.015 2.012 0.011 0.905
$\alpha_{2}$ 0.042 0.466 0.033 0.310 -0.001 0.227 -0.006 0.100
$\alpha_{3}$ 0.000 0.222 0.004 0.152 0.000 0.099 -0.003 0.045
$\gamma_{0}$0.233 1.612 -0.067 1.020 -0.076 0.703 -0.004 0.308
$\gamma_{1}$ 0.406 4.332 -0.024 3.174 -0.178 2.585 -0.050 1.110
$\gamma_{2}$ 0.025 0.231 0.008 0.154 0.010 0.103 0.000 0.045
$\gamma_{3}$ 0.034 0.233 0.010 0.146 0.009 0.103 0.002 0.045
$\theta$ 0.014 0.380 0.018 0.311 0.035 0.260 0.009 0.113
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.138 2.592 0.142 1.671 0.004 0.936 0.000 0.410
$\alpha_{1}$0.986 11.090 -0.665 6.325 -0.174 4.042 0.003 1.788
$\alpha_{2}$ 0.065 1.255 0.112 0.756 0.023 0.447 0.000 0.207
$\alpha_{3}$0.006 0.589 0.019 0.345 0.003 0.224 0.003 0.092
$\gamma_{0}$0.123 3.350 -0.298 2.455 -0.081 1.379 -0.047 0.628
$\gamma_{1}$ 0.635 7.424 0.141 5.668 -0.081 4.255 -0.040 2.159
$\gamma_{2}$ 0.016 0.484 0.037 0.349 0.009 0.205 0.005 0.093
$\gamma_{3}$ 0.009 0.517 0.026 0.336 0.000 0.208 0.007 0.092
$\theta$0.029 0.446 0.025 0.414 0.034 0.381 0.016 0.220
Sample size ($T$) 50 100 200 1000
Note: The data generating process follows Matsumura and Otani (2023b).

### v2

@@
 Abstract
-We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in a log-linear model with homogeneous goods markets.
-First, we derive such conditions.
-Second, Monte Carlo simulations confirm that in a log-linear model, incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters.
-Keywords: Conduct parameters, Homogenous Goods Market, Mathematical Programming with Equilibrium Constraints, Monte Carlo simulation
+We propose a constrained generalized method of moments (GMM) estimator with some equilibrium uniqueness conditions for estimating the conduct parameter in a log-linear model with homogeneous goods markets.
+Monte Carlo simulations demonstrate that merely imposing parameter restrictions leads to not just inaccurate estimations but also some numerical issues, and adding the equilibrium uniqueness conditions resolves them.
+We also suggest a formulation of the GMM estimation to further avoid the numerical issues.
+Keywords: Conduct parameters, Homogeneous Goods Market, Mathematical Programming with Equilibrium Constraints, Monte Carlo simulation
 JEL Codes: C5, C13, L1
 1 Introduction
-Measuring competitiveness is a crucial task in the empirical industrial organization literature.
-Conduct parameter is considered a useful measure of competitiveness.
-However, it cannot be directly measured from data because data usually lack information about marginal cost.
-Therefore, researchers aim to identify and estimate the conduct parameter.
-As the simplest specification, Bresnahan (1982) considers identification of conduct parameters for the linear model.
-Matsumura and Otani (2023b, a) resolves the conflict on some identification problems between Bresnahan (1982) and Perloff and Shen (2012).
-However, researchers often implement alternative specifications, such as the log-linear model (Okazaki et al.
-2022, Mérel 2009).
-In the context of the log-linear model, these papers identify estimation issues, with some estimated conduct parameters being unrealistically low or even negative.
-This raises doubts about the methodology, despite the identification strategy outlined by Lau (1982).
-This is an obstacle to choosing a better specification of the demand and supply functions.
-To overcome the problem, we propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices as constraints.
-First, we prove that a unique equilibrium exists under certain theoretical conditions.
-As far as we know, this is a new result.
-Second, we show that incorporating equilibrium existence and uniqueness conditions resolves the above estimation problems and makes the model work in estimating conduct parameters even in the log-linear model.
-Hence, our results support those of Lau (1982) numerically.
+Measuring competitiveness is central in empirical IO, and the conduct parameter is a widely used proxy.
+Since marginal cost is typically unobserved, researchers identify and estimate the conduct parameter indirectly.
+bresnahan1982oligopoly develops an identification strategy under linear demand and marginal cost, and matsumura2023resolving provide detailed conditions for this case.
+Yet, empirical work often employs log-linear models (okazaki2022excess, merel2009measuring), which tend to produce implausibly low or negative estimates, even though identification conditions for general models are available from lau1982identifying and matsumura2025revisiting.
+This casts doubt on the methodology and complicates model selection.
+We address this by proposing a constrained GMM estimator that incorporates theoretical conditions for the uniqueness of equilibrium.
+First, we derive new conditions guaranteeing a unique equilibrium.
+Second, Monte Carlo simulations show that parameter restrictions alone yield inaccurate estimates and numerical errors, while adding equilibrium conditions resolves them.
+We also propose a modified GMM formulation that further mitigates these issues.
 2 Model
 Consider data with $T$ markets with homogeneous products. Assume there are $N$ firms in each market. Let $t=1,\ldots,T$ be the index of markets. Then, we obtain the supply equation:
-$P_{t}=-\theta\frac{\partial P_{t}(Q_{t})}{\partial Q_{t}}Q_{t}+MC_{t}(Q_{t}),$ (1)
-where $Q_{t}$ is the aggregate quantity, $P_{t}(Q_{t})$ is the demand function, $MC_{t}(Q_{t})$ is the marginal cost function, and $\theta\in[0,1]$, which is the conduct parameter.
-The equation nests perfect competition $(\theta=0)$, Cournot competition $(\theta=1/N)$, and perfect collusion $(\theta=$ 1)1)1 ) (Bresnahan 1982).
-Consider an econometric model. Assume that the demand and the marginal cost functions are written as follows:
+$P_{t}+\theta Q_{t}P_{t}^{\prime}(Q_{t})=MC_{t}(Q_{t}),$ (1)
+where $Q_{t}$ is the aggregate quantity, $P_{t}(Q_{t})$ is the inverse demand function, $MC_{t}(Q_{t})$ is the marginal cost function, and $\theta\in[0,1]$ is the conduct parameter.
+The equation nests perfect competition $(\theta=0)$, Cournot competition $(\theta=1/N)$, and perfect collusion $(\theta=$ $1)$.
+See bresnahan1982oligopoly for the details.
+Consider an econometric model. Assume that the inverse demand and the marginal cost functions are given as
 $P_{t}=f(Q_{t},X^{d}_{t},\varepsilon^{d}_{t},\alpha),$
@@
 where $X^{d}_{t}$ and $X^{c}_{t}$ are the vector of exogenous variables, $\varepsilon^{d}_{t}$ and $\varepsilon^{c}_{t}$ are the error terms, and $\alpha$ and $\gamma$ are the vector of parameters.
-We also have the demand- and supply-side instruments, $Z^{d}_{t}$ and $Z^{c}_{t}$, and assume that the error terms satisfy the mean independence condition $E[\varepsilon^{d}_{t}\mid X^{d}_{t},Z^{d}_{t}]=E[\varepsilon^{c}_{t}\mid X^{c}_{t},Z^{c}_{t}]=0$.
+We allow $X^{d}_{t}$ and $X^{c}_{t}$ to have common variables, but assume that there is at least one demand variable and one cost variable that are mutually excluded.
+We also have the demand- and supply-side instruments, $Z^{d}_{t}$ and $Z^{c}_{t}$, and assume that the error terms satisfy the mean independence condition, $E[\varepsilon^{d}_{t}\mid X^{d}_{t},Z^{d}_{t}]=E[\varepsilon^{c}_{t}\mid X^{c}_{t},Z^{c}_{t}]=0$.
+The identification of the conduct parameter is indirectly characterized by lau1982identifying:
+Theorem 1.
+Under the assumption that the industry inverse demand and cost functions are twice continuously differentiable, the index of competitiveness $\theta$ cannot be identified from data on industry price and output and other exogenous variables alone if and only if the industry inverse demand function is separable in $X^{d}$, that is, $f(Q,r(X^{d}))$, but not take the form $P=Q^{-1/\theta}r(X^{d})+s(Q)$.
+This theorem implies that the conduct parameter is identified if the inverse demand function is not separable.
+A demand rotation instrument (bresnahan1982oligopoly) achieves this.
+See Appendix A.1 for the details of the definition of separability.
 2.1 Log-linear demand and log-linear marginal cost
-Consider a log-linear model, which is a typical specification. The demand and marginal cost functions are specified as
+Consider a log-linear model, which is a typical specification. The inverse demand and marginal cost functions are specified as
 $\log P_{t}$ $=\alpha_{0}-(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t},$ (2)
 $\log MC_{t}$ $=\gamma_{0}+\gamma_{1}\log Q_{t}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t},$ (3)
-where $Y_{t}$ is an excluded demand shifter, $W_{t}$ and $R_{t}$ are excluded cost shifters, and $Z_{t}^{R}$ is Bresnahan’s demand rotation instrument. Then, Equation (1) is written as
+where $Y_{t}$ and $Z_{t}^{R}$ are excluded demand shifters and $W_{t}$ and $R_{t}$ are excluded cost shifters.
+When $Y_{t}$ and $Z_{t}^{R}$ vary without changing the equilibrium quantity, they work as the demand rotation instrument.
+Then, (1) is written as
 $P_{t}$ $=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+MC_{t}.$ (4)
-By taking logarithm of Equation (4) and substituting Equation (3), we obtain
+By taking logarithm of (4) and substituting (3), we obtain
 $\log P_{t}=-\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))+\gamma_{0}+\gamma_{1}\log Q_{t}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}.$ (5)
-In this model, the number of equilibrium prices varies.
-Although it is a widely known model, to our knowledge, there is no paper examining this.
-Thus, we derive the conditions for the uniqueness in the following proposition which are used in our estimation.
+The intersection of (2) and (5) determines the equilibrium, but there could be multiple equilibria.
+Although this model is widely known, no paper has examined the multiple equilibria problem to our knowledge.
+The next proposition provides the conditions for uniqueness.
+The proof is in the online appendix A.2.
 Proposition 1.
-Assume that $\alpha_{1}+\alpha_{2}Z^{R}_{t}\neq 0$.
-Let $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y_{t}+\varepsilon^{d}_{t}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}+\gamma_{2}\log W_{t}+\gamma_{3}\log R_{t}+\varepsilon^{c}_{t}$.
-The number of equilibrium prices $P_{t}^{*}>0$ is determined as follows:
-  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, there is no equilibrium price,
-  • When $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0,
-    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\neq 1$, there is a unique equilibrium price,
-    • If $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, there are infinitely many equilibrium prices when $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, but there is no equilibrium price otherwise.
-See the online appendix A.1 for the proof.
+Assume that $\alpha_{1}+\alpha_{2}Z^{R}\neq 0$.
+Let $\Xi=\gamma_{0}+\gamma_{1}\frac{\alpha_{0}+\alpha_{3}\log Y+\varepsilon^{d}}{\alpha_{1}+\alpha_{2}Z^{R}}+\gamma_{2}\log W+\gamma_{3}\log R+\varepsilon^{c}$.
+The number of equilibria is determined as follows:
+  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R})\leq 0$, there is no equilibrium,
+  • When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R})>0$,
+    • If $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}\neq 0$, there is a unique equilibrium,
+    • If $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}=0$, there are infinitely many equilibria when $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, but there is no equilibrium otherwise.
+The condition $1-\theta(\alpha_{1}+\alpha_{2}Z^{R})>0$ rules out the region where the log transformation leaves its domain, which corresponds to implausibly elastic demand combined with large conduct.
+The assumption $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}\neq 0$ excludes the knife-edge case in which the (pseudo) supply and demand are exactly parallel so that every price could be an equilibrium; both restrictions are technical and do not bind in regular empirical settings.
 3 Estimation
-To estimate parameters in the demand and supply equations, we use GMM estimation. Among GMM estimators, we apply the nonlinear system two-stage-least-squares (N2SLS) using Equation (2) and (5).
-Let $\xi=(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\theta)$ be the vector of parameters.
-Given the demand equation (2) and the supply equation (5), we can write the error terms in the demand and supply equation as
+Let $\xi=(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\theta)$ be the vector of the parameters in the model.
+We use the GMM for the estimation.
+Among GMM estimators, we apply the nonlinear system two-stage-least-squares (N2SLS) using (2) and (5).
+We rewrite the demand equation (2) and the supply equation (5) as
 ${\varepsilon}_{t}^{d}(\xi)$ $=\log P_{t}-\alpha_{0}+(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}-\alpha_{3}\log Y_{t},$ (6)
 ${\varepsilon}_{t}^{c}(\xi)$ $=\log P_{t}+\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))-\gamma_{0}-\gamma_{1}\log Q_{t}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t}.$ (7)
-To estimate the parameters, we convert the conditional moment conditions, $E[\varepsilon_{t}^{d}\mid Z_{t}^{d}]=E[\varepsilon_{t}^{c}\mid Z_{t}^{c}]=0$, into unconditional moment conditions, $E[\varepsilon_{t}^{d}Z_{t}^{d}]=E[\varepsilon_{t}^{c}Z_{t}^{c}]=0$.
-Using Equations (6) and (7), we construct the sample analog of the moment conditions:
+To estimate the parameters, we convert the conditional moments, $E[\varepsilon_{t}^{d}\mid Z_{t}^{d}]=E[\varepsilon_{t}^{c}\mid Z_{t}^{c}]=0$, into unconditional moments, $E[\varepsilon_{t}^{d}Z_{t}^{d}]=E[\varepsilon_{t}^{c}Z_{t}^{c}]=0$.
+Using Equations (6) and (7), we construct the sample analog of the unconditional moments:
 $g(\xi)=\left[\begin{array}[]{l}\frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{d}_{t}(\xi)Z_{t}^{d}\\ \frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{c}_{t}(\xi)Z_{t}^{c}\end{array}\right].$
-We define the GMM estimator as the vector $\xi^{*}$ that solves the problem,
-$\min_{\xi}\ g(\xi)^{\top}Wg(\xi)$ (8)
+We define the N2SLS estimator as the solution to the problem,
+$\xi^{*}=\arg\min_{\xi}\ g(\xi)^{\top}Wg(\xi)$ (8)
 where the weight matrix $W$ is defined as
 $W=\left[\frac{1}{T}\sum_{t=1}^{T}Z_{t}^{\top}Z_{t}\right]^{-1}\text{ where }Z_{t}=\left[\begin{array}[]{ll}Z_{t}^{d\top}&0\\ 0&Z_{t}^{c\top}\end{array}\right].$ (11)
-The solution to (8) is called the N2SLS estimator (Wooldridge 2010).
-Using Proposition 1, we impose the following constraints:
+We also add the following constraints based on Proposition 1 to (8):
 $0\leq\theta\leq 1,$ (12)
@@
 Constraint (12) is a standard assumption on the conduct parameter.
-Constraints (13) and (14) relate to the uniqueness of equilibrium prices.
-Constraint (13) implies the downward-sloping demand and upward-sloping marginal cost.
-See online appendix A.2 for the detailed setting.
+Constraint (13) implies the downward-sloping demand and upward-sloping marginal cost, which guarantees that $\gamma_{1}+\alpha_{1}+\alpha_{2}Z^{R}\neq 0$.
+Constraint (14) relates to the uniqueness of equilibrium.
+See the detailed simulation setting in the online appendix A.3.
 4 Simulation results
-We compare N2SLS estimation with and without Constraints (12), (13), and (14).
-Table 1 presents the results.
-First, as reported in the literature (Mérel 2009, Okazaki et al.
-2022), N2SLS without constraints converges to implausibly low and negative conduct parameters.
-This is because the algorithm searches the area in which no equilibrium price and quantity exist and then finds unreasonable local optima that might have a lower objective function than at the true values.
-Second, N2SLS with constraints converges to the reasonable values.
-As the sample size increases, the bias and root-mean-square error (RMSE) decrease to levels comparable to linear models.
-Third, we find an ad hoc method that uses Equation (3) to compute the residual in the supply estimation, $\varepsilon_{t}^{c}$, and Equation (4) as an equality constraint with Constraints (12), (13), and (14).
-Then, the bias and RMSE of conduct parameter $\theta$ are reduced to 0.014 and 0.217 in Panel (c), although the results do not dominate the results in Panel (b) for all parameters.
-Therefore, incorporating equilibrium existence and uniqueness conditions simply is helpful for conduct parameter estimation in nonlinear models such as a log-linear model.
-See online appendix A.4 for additional experiments under different variances of errors and the results of a linear model.
+We compare N2SLS estimations with and without constraints in Table 1.
+Panel (a) shows that, without constraints, the estimator fails to recover $\gamma_{0}$ and $\theta$, replicating known issues due to the flat objective function and invalid search regions without equilibrium (Appendix A.4).
+Panel (b), which imposes Constraint (12), improves estimation in large samples via the domain restriction.
+[^1] Constraints (13) and (14) alone yield severe bias; see Table 9 and Appendix A.5.
+However, in small samples, demand parameter estimates degrade and convergence declines.
+When convergence fails, $\alpha_{1}$ becomes large, rendering $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})<0$ and causing numerical errors inside the log term in (1).
+Adding constraints (13) and (14) in Panel (c) improves small-sample convergence and demand accuracy, though convergence is not guaranteed.
+In large samples, performance surpasses that of Panel (b) for some parameters.
+To address convergence failure, we propose an alternative formulation (Table 2) that computes $\varepsilon^{c}_{t}$ via (3) and enforces Equation (4) as a constraint, along with Constraints (12)–(14).
+[^2] See Appendix A.3 for details.
+This avoids log terms in both objective and constraints, achieving 100% convergence and reducing $\theta$’s bias and RMSE to 0.014 and 0.217, though not dominating Panel (c) across all parameters.
+In sum, incorporating equilibrium uniqueness conditions and eliminating log terms greatly improves conduct parameter estimation.
+Additional experiments appear in Appendix A.5.
 Table 1: Performance comparison
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$1.070 7.012 -0.021 5.110 0.365 2.207 0.400 2.030
-$\alpha_{1}$0.164 1.060 -0.001 0.782 0.073 0.458 0.096 0.574
-$\alpha_{2}$0.011 0.104 -0.006 0.071 0.002 0.033 0.005 0.043
-$\alpha_{3}$0.101 0.619 -0.005 0.474 0.021 0.198 0.029 0.187
-$\gamma_{0}$ 9.735 15.743 9.636 10.870 13.173 13.269 13.294 13.351
-$\gamma_{1}$0.070 1.624 -0.177 0.469 -0.184 0.248 -0.177 0.220
-$\gamma_{2}$0.034 0.939 -0.098 0.317 -0.090 0.152 -0.080 0.127
-$\gamma_{3}$0.047 0.750 -0.091 0.311 -0.098 0.156 -0.085 0.133
-$\theta$3e+05 3e+06 -2e+05 2e+06 -8e+04 9e+04 -9e+04 1e+05
-Runs converged (%) 99.500 99.800 98.600 98.400
-Sample size ($T$) 100 200 1000 1500
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.905 6.954 0.120 5.001 0.072 2.042 0.052 1.563
-$\alpha_{1}$0.141 1.053 0.018 0.768 0.010 0.313 0.008 0.241
-$\alpha_{2}$0.006 0.101 0.000 0.062 0.001 0.024 0.001 0.019
-$\alpha_{3}$0.088 0.620 0.007 0.475 -0.001 0.186 0.003 0.152
-$\gamma_{0}$1.748 14.206 -0.938 6.428 0.015 1.995 0.163 1.570
-$\gamma_{1}$ 0.254 1.927 0.129 0.825 0.018 0.226 0.003 0.170
-$\gamma_{2}$ 0.117 1.083 0.049 0.467 0.008 0.148 0.007 0.116
-$\gamma_{3}$ 0.098 0.890 0.058 0.478 -0.001 0.148 0.003 0.118
-$\theta$0.100 0.441 -0.072 0.424 -0.121 0.351 -0.148 0.333
-Runs converged (%) 99.600 99.900 100.000 100.000
-Sample size ($T$) 100 200 1000 1500
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
-$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
-$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
-$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
-$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
-$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
-$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
-$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
-$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
-Runs converged (%) 100.000 100.000 100.000 100.000
-Sample size ($T$) 100 200 1000 1500
+(a) N2SLS without Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.070 7.012 -0.021 5.110 0.365 2.207 0.400 2.030 $\alpha_{1}$ -0.164 1.060 -0.001 0.782 0.073 0.458 0.096 0.574 $\alpha_{2}$ -0.011 0.104 -0.006 0.071 0.002 0.033 0.005 0.043 $\alpha_{3}$ -0.101 0.619 -0.005 0.474 0.021 0.198 0.029 0.187 $\gamma_{0}$ 9.735 15.743 9.636 10.870 13.173 13.269 13.294 13.351 $\gamma_{1}$ -0.070 1.624 -0.177 0.469 -0.184 0.248 -0.177 0.220 $\gamma_{2}$ -0.034 0.939 -0.098 0.317 -0.090 0.152 -0.080 0.127 $\gamma_{3}$ -0.047 0.750 -0.091 0.311 -0.098 0.156 -0.085 0.133 $\theta$ -3e+05 3e+06 -2e+05 2e+06 -8e+04 9e+04 -9e+04 1e+05 Runs converged (%) 99.500 99.800 98.600 98.400 Sample size ($T$) 100 200 1000 1500
+(b) N2SLS with Constraints (12)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.922 8.603 -0.068 5.116 0.037 2.035 0.000 1.556 $\alpha_{1}$ -0.299 1.314 -0.010 0.785 0.005 0.312 0.000 0.240 $\alpha_{2}$ -0.013 0.104 -0.002 0.063 0.001 0.024 0.000 0.019 $\alpha_{3}$ -0.165 0.774 -0.007 0.472 -0.004 0.185 -0.001 0.152 $\gamma_{0}$ -1.767 14.394 -1.001 6.530 -0.208 1.993 -0.156 1.566 $\gamma_{1}$ 0.255 1.949 0.132 0.838 0.034 0.229 0.027 0.174 $\gamma_{2}$ 0.125 1.097 0.053 0.475 0.017 0.150 0.019 0.119 $\gamma_{3}$ 0.099 0.903 0.062 0.481 0.007 0.149 0.014 0.120 $\theta$ -0.098 0.441 -0.060 0.421 -0.061 0.319 -0.058 0.281 Runs converged (%) 98.100 98.700 100.000 100.000 Sample size ($T$) 100 200 1000 1500
+(c) N2SLS with Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.905 6.954 0.120 5.001 0.072 2.042 0.052 1.563 $\alpha_{1}$ -0.141 1.053 0.018 0.768 0.010 0.313 0.008 0.241 $\alpha_{2}$ -0.006 0.101 0.000 0.062 0.001 0.024 0.001 0.019 $\alpha_{3}$ -0.088 0.620 0.007 0.475 -0.001 0.186 0.003 0.152 $\gamma_{0}$ -1.748 14.206 -0.938 6.428 0.015 1.995 0.163 1.570 $\gamma_{1}$ 0.254 1.927 0.129 0.825 0.018 0.226 0.003 0.170 $\gamma_{2}$ 0.117 1.083 0.049 0.467 0.008 0.148 0.007 0.116 $\gamma_{3}$ 0.098 0.890 0.058 0.478 -0.001 0.148 0.003 0.118 $\theta$ -0.100 0.441 -0.072 0.424 -0.121 0.351 -0.148 0.333 Runs converged (%) 99.600 99.900 100.000 100.000 Sample size ($T$) 100 200 1000 1500
 Note: The error terms are drawn from a normal distribution, $N(0,\sigma)$.
 True values: $\alpha_{0}=20.0,\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=1.0,\gamma_{0}=5.0,\gamma_{1}=1.0,\gamma_{2}=1.0,\gamma_{3}=1.0,\theta=0.5$ and $\sigma=1.0$.
-See online appendix A.2 for the setting.
-5 Conclusion
-We propose a constrained generalized method of moments estimator (GMM) incorporating theoretical conditions for the unique existence of equilibrium prices for estimating conduct parameters in homogeneous goods markets.
-First, we derive the conditions.
-Second, Monte Carlo simulations confirm that incorporating the conditions resolves the problems of implausibly low or negative values of conduct parameters in a log-linear model, as reported in the literature.
+See online appendix A.3 and matsumura2024challenges for the setting.
+Table 2: Ad hoc method using (3) to compute $\varepsilon_{t}^{c}$ and (4) with Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555 $\alpha_{1}$ -0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240 $\alpha_{2}$ -0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020 $\alpha_{3}$ -0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152 $\gamma_{0}$ -5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533 $\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175 $\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120 $\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121 $\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217 Runs converged (%) 100.000 100.000 100.000 100.000 Sample size ($T$) 100 200 1000 1500
+5 Discussion
+Two concerns surround the conduct parameter approach: the difficulty of interpreting intermediate or extreme values, and the critique by corts1999conduct that it may understate market power under collusion.
+[^3] As magnolfi2022comparison note, this critique does not apply when the data stem from a static model.
+We show that implausible estimates in log-linear models often stem from numerical issues—especially when equilibrium conditions are omitted—rather than conceptual flaws.
+Addressing these issues yields more stable and interpretable results.
+This distinction matters: misattributing numerical artifacts to theoretical limits risks dismissing a useful tool.
+Our findings aim to encourage more constructive use of the conduct parameter approach.
 Acknowledgments
-We thank Jeremy Fox, Yelda Gungor, and Isabelle Perrigne for valuable comments.
-This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
+We thank Jeremy Fox, Yelda Gungor, and Isabelle Perrigne for their valuable comments.
+We were supported by JST ERATO Grant Number JPMJER2301 and Otani was supported by JSPS Grant-in-Aid (KAKENHI) for Young Researcher JSPS 24K22604 and 25K16620.
 Appendix A Online appendix
-A.1 Existence and uniqueness of equilibrium prices
+A.1 Summary of Goldman and Uzawa (1964)
+goldmanNote1964 investigate separability concepts in demand analysis.
+Let $N$ be a set of variables, $N=\{1,\ldots,n\}$.
+Consider a partition of $N$ into $S$ parts, $\{N^{1},\ldots,N^{S}\}$ such that $N=\bigcup_{s=1}^{S}N^{s}$ and $N^{s}\cap N^{t}=\emptyset$.
+A function is weakly separable with respect to a partition if
+$\frac{\partial}{\partial x_{k}}\left(\frac{\partial f(x)/\partial x_{i}}{\partial f(x)/\partial x_{j}}\right)=0,\quad i,j\in N^{s},k\notin N^{s}.$ (15)
+This implies that, for example, the marginal rate of substitution between $i$ and $j$ in the same partition is independent of the quantities of commodities outside $N^{s}$.
+In Theorem 2, goldmanNote1964 characterizes the function that satisfies the weak separability.
+Theorem 2 (Theorem 2 in goldmanNote1964).
+A function $f(x)$ is weakly separable with respect to a partition $\{N^{1},...,N^{s}\}$ if, and only if, $f(x)$ is of the form:
+$f(X)=\Phi(f^{1}(x^{(1)}),\ldots,f^{s}(x^{(s)}))$ (16)
+where $\Phi(f^{1},\ldots,f^{s})$ is a function of $S$ variables and, for each $s$, $f^{s}(x^{(s)})$ is a function subvector $x^{(s)}$ alone.
+Let $\{Q,X_{1}^{d},\ldots,X_{K}^{d}\}$ be the set of dependent variables and divide the variables into two groups $(Q,X^{d})$.
+Then, as the inverse demand function $P(Q,X^{d})=f(Q,r(X^{d}))$ in Theorem 1 has the form in the above theorem, we can say that the conduct parameter is not identified when the inverse demand function is separable.
+In contrast, we can see that the linear demand function in bresnahan1982oligopoly and the log-linear demand function in this paper are not separable. For the linear case,
+$\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}=\frac{\alpha_{2}Q}{\alpha_{3}}\Longrightarrow\frac{\partial}{\partial Q}\left(\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}\right)=\frac{\alpha_{2}}{\alpha_{3}}\neq 0.$ (17)
+The last part, $\frac{\alpha_{2}}{\alpha_{3}}\neq 0$, is concluded from the identification condition in matsumura2023resolving. For the log-linear case, we have
+$\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}=\frac{\alpha_{2}Y\log Q}{\alpha_{3}}\Longrightarrow\frac{\partial}{\partial Q}\left(\frac{\partial P(Q,X^{d})/\partial Z^{R}}{\partial P(Q,X^{d})/\partial Y}\right)=\frac{\alpha_{2}Y}{\alpha_{3}Q}\neq 0.$ (18)
+Thus, again when $\frac{\alpha_{2}}{\alpha_{3}}$ is nonzero, which emphasizes the role of the demand shifters, the inverse demand function is not separable.
+Note that due to the log specification, $Y$ and $Q$ are positive.
+A.2 Existence and uniqueness of equilibrium prices
 Proposition 1 proposes the conditions for the unique existence of $P_{t}(>0)$ solving the demand equation (2) and supply equation (5) for $P_{t}$ under $\theta\in[0,1]$.
@@
 Thus $P_{t}^{*}=0$ holds.
-However, our interest is a unique positive equilibrium price, so we seek conditions in which there is a unique positive price that satisfies $\Delta(P_{t})>0$.
+However, our interest is a unique positive equilibrium price, so we seek conditions in which there is a unique positive price that satisfies $\Delta(P_{t})=0$.
 When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, (I) is always negative on $P_{t}>0$.
-In contrast, (II) is non-negative regardless of the sign of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ on $P_{t}>0$.
+In contrast, (II) is non-negative regardless of the sign of $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ on $P_{t}>0$.
 Therefore, $\Delta(P_{t})$ is always negative on $P_{t}>0$, which implies that there is no positive equilibrium price.
@@
 $\frac{d^{2}}{dP_{t}^{2}}\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$ $=-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}\left(-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1\right)\exp(\Xi)P_{t}^{-\frac{\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-2}.$ (22)
-Therefore, the shape of (II) on $P_{t}>0$ is changed based on the value of $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ and the number that (I) and (II) cross on $P_{t}>0$ is also determined.
-The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})<0$ is illustrated in Figure 1(a).
+Therefore, the shape of (II) on $P_{t}>0$ is changed based on the value of $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})$ and the number that (I) and (II) cross on $P_{t}>0$ is also determined.
+The case $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})<0$ is illustrated in Figure 1(a).
 In this case, (II) is a monotone decreasing convex function because the first and second derivatives are both negative.
 As (I) is monotone increasing in $P_{t}>0$, (I) and (II) cross only once on $P_{t}>0$.
-The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\in[0,1)$ is illustrated in Figure 1(b).
+The case $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})\in[0,1)$ is illustrated in Figure 1(b).
 In this case, (II) becomes a monotone increasing concave function passing through the origin because the first derivative is positive and the second derivative is negative.
-As the value of the first derivative is infinity at $P_{t}=0$ since $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})-1<0$, (II) should be greater than (I) around $P_{t}=0$.
-Then as $P_{t}$ becomes large starting from zero, the difference in both terms becomes small, and both terms cross only once on $P_{t}>0$ eventually.
-When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, (16) becomes
+As the value of the first derivative is infinity at $P_{t}=0$ since $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})-1<0$, (II) should be greater than (I) around $P_{t}=0$.
+Then, as $P_{t}$ becomes large starting from zero, the difference in both terms becomes small, and both terms cross only once on $P_{t}>0$ eventually.
+When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})=1$, (20) becomes
 $0=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)]P_{t}.$
 If $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})=\exp(\Xi)$, the above equation holds for any $P_{t}$.
-Thus there are infinitely many equilibrium prices on $P_{t}>0$, which is illustrated in Figure 1(c).
+Thus, there are infinitely many equilibrium prices on $P_{t}>0$, which is illustrated in Figure 1(c).
 When $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq\exp(\Xi)$, there is no equilibrium price on $P_{t}>0$ because the right-hand side of the equation is always non-zero on $P_{t}>0$, which is illustrated in Figure 1(d).
-The case $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})>1$ is illustrated in Figure 1(e).
+The case $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R})>1$ is illustrated in Figure 1(e).
 In this case, (II) becomes a monotone increasing convex function passing through the origin because the first and second derivatives are positive.
@@
 Figure 1: The illustrations of the equilibrium existence and uniqueness conditions
-$P_{t}$$\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$$[1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})]P_{t}$
+$P_{t}$$\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$$1-\theta(\alpha_{1}+\alpha_{2}Z_{t}^{R})]P_{t}$
 (a) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}<0$
@@
 (e) $\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}>1$
-Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0.
+Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>0$.
 The thick line represents the first term and the dotted line represents the second term in (20).
 From $\Delta(P_{t})=0$, the positive equilibrium price can be written as
-00 0 $=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$
-00 0 $=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$
+$0$ $=[1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})]P_{t}-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}}$
+$0$ $=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})-\exp(\Xi)P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$
 $P_{t}^{\frac{-\gamma_{1}}{\alpha_{1}+\alpha_{2}Z^{R}_{t}}-1}$ $=\frac{1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})}{\exp(\Xi)}$
@@
 Thus, there should not be any equilibrium.
-Hereafter, assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0.
-When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand equation and supply equation coincides.
+Hereafter, assume that $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>0$.
+When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand equation and supply equation coincides.
 Thus there are infinitely many equilibria in the model.
-When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand and supply equations have a same slope and different intercepts, which means that both equations become parallel.
+When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$, the demand and supply equations have a same slope and different intercepts, which means that both equations become parallel.
 Thus, there is no equilibrium.
-When $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, the demand and supply equations have different slopes, and hence we can find a unique equilibrium.
+When $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, the demand and supply equations have different slopes, and hence we can find a unique equilibrium.
 Figure 2: The illustrations of the equilibrium existence and uniqueness conditions
-[…]
-(a) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
-[…]
-(b) $-\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
-[…]
+$Q_{t}$$P_{t}$SSDD
+(a) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)=1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
+$Q_{t}$$P_{t}$SSDD
+(b) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})=1$ and $\exp(\Xi)\neq 1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})$
+$Q_{t}$$P_{t}$SSDD
 (c) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$, and $\gamma_{1}>0$
-[…]
+$Q_{t}$$P_{t}$SSDD
 (d) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}\leq 0$, and $\gamma_{1}>0$
-[…]
+$Q_{t}$$P_{t}$SSDD
 (e) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$, and $\gamma_{1}\leq 0$
-[…]
+$Q_{t}$$P_{t}$DDSS
 (f) $\gamma_{1}/(\alpha_{1}+\alpha_{2}Z^{R}_{t})\neq 1$, $\alpha_{1}+\alpha_{2}Z^{R}_{t}\leq 0$, and $\gamma_{1}\leq 0$
-Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $01-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>$ 0.
+Note: As there is no equilibrium when $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})\leq 0$, the above figures assume that $1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})>0$.
 Case (a) and (b) also allow both the demand and supply equations to be downward sloping.
-A.2 Simulation and estimation procedure
+A.3 Simulation and estimation procedure
 To generate the simulation data, for each model, we first generate the exogenous variables $Y_{t},Z^{R}_{t},W_{t},R_{t},H_{t}$, and $K_{t}$ and the error terms $\varepsilon_{t}^{c}$ and $\varepsilon_{t}^{d}$ based on the data generation process in Table 3.
@@
 We generate 1000 data sets of 100, 200, 1000, 1500 markets.
-We jointly estimate the demand and supply parameters by the simultaneous equation model (Wooldridge 2010) from the true values.
+We jointly estimate the demand and supply parameters by the simultaneous equation model (wooldridge2010econometric) from the true values.
 We use state-of-the-art constrained optimization solvers, i.e., Ipopt.jl which implements an interior point line search filter method that aims to find a local solution of nonlinear programming problems.
@@
 Note: $\sigma=\{0.5,1.0,2.0\}$. $N:$ Normal distribution. $U:$ Uniform distribution.
-A.3 Why does the estimation without equilibrium conditions converge to the extremely low conduct parameters?
-We should interpret reasons why the estimation without equilibrium conditions converges to the negative conduct parameters, as shown in Panel (a) in Table 1, although the obtained parameters cannot be interpreted as the exogenous elements determining equilibrium outcomes.
-First, we rewrite the terms in Equation (7) into
-$\log(1-\theta C_{t})-\gamma_{0}$
-where $C_{t}=\alpha_{1}+\alpha_{2}Z^{R}_{t}$ is given in the demand estimation.
-If the model allows $\theta<0$, then $1-\theta C_{t}$ can take an arbitrarily large value through log-transformation relative to $\gamma_{0}$.
-For example, if $0001-\theta C_{t}=1,000,$ 000, then $\log(1-\theta C_{t})=13.8$, which has a small contribution relative to $\gamma_{0}$ for evaluating the moment conditions.
-A.4 Additional experiments
-Additional results for different $\sigma$ are shown in Tables 3 and 4. As a summary, the main findings in the main text are robust.
-In Tables 6 and 7, we illustrate that N2SLS with Constraints (12) works for the linear model as in Matsumura and Otani (2023b).
+The ad-hoc estimation for Table 2 is formulated as follows. Let $\Phi=(\xi,\{MC_{t}\}_{t=1,\ldots,T})$. Consider a constrained minimization problem such that
+$\Phi^{*}=\arg\min_{\Phi}\ g(\Phi)^{\top}Wg(\Phi),$ (25)
+subject to
+$(\Phi)=\left[\begin{array}[]{l}\frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{d}_{t}(\Phi)Z_{t}^{d}\\ \frac{1}{T}\sum_{t=1}^{T}{\varepsilon}^{c}_{t}(\Phi)Z_{t}^{c}\end{array}\right],$ (29)
+${\varepsilon}_{t}^{d}(\Phi)=\log P_{t}-\alpha_{0}+(\alpha_{1}+\alpha_{2}Z^{R}_{t})\log Q_{t}-\alpha_{3}\log Y_{t},$ (30)
+${\varepsilon}_{t}^{c}(\Phi)=\log MC_{t}-\gamma_{0}-\gamma_{1}\log Q_{t}-\gamma_{2}\log W_{t}-\gamma_{3}\log R_{t},$ (31)
+$P_{t}=\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t})P_{t}+MC_{t},$ (32)
+$MC_{t}\geq 0,$ (33)
+where the matrix $W$ is the same as the original N2SLS.
+A.4 Why does the estimation without equilibrium conditions converge to the extremely low conduct parameters?
+In the literature, it is often observed that estimation without imposing equilibrium conditions tends to converge to negative values of the conduct parameter.
+We confirm this pattern in our Monte Carlo simulations, particularly in Panel (a) of Table 1.
+To understand this phenomenon, consider the supply equation (5), which includes the following term:
+$-\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}{t}))+\gamma{0}.$ (34)
+If the model allows for $\theta<0$, the term inside the logarithm can become arbitrarily large in magnitude by choosing a sufficiently negative $\theta$ when $\alpha_{1}+\alpha_{2}Z^{R}_{t}>0$:
+$\lim_{\theta\rightarrow-\infty}\log(1-\theta(\alpha_{1}+\alpha_{2}Z^{R}_{t}))\rightarrow\infty.$ (35)
+However, since $\theta$ appears inside a logarithmic function, changes in $\theta$—even when large—have a relatively small impact on the GMM objective function.
+Moreover, the effect of changes in $\theta$ can be offset by adjusting $\gamma_{0}$, with only modest changes in its value.
+This implies that the GMM objective function can be very flat in the neighborhood of the true values of $\theta$ and $\gamma_{0}$.
+We verify this implication by examining the shape of the GMM objective function.
+Figure 3 plots the objective function on the $\theta$$\gamma_{0}$ plane, fixing all other parameters at their true values and using one simulated dataset.
+The same pattern is observed across other simulation datasets as well.
+As the figure clearly shows, the GMM objective function is nearly flat around the true value of $\gamma_{0}$, and it allows for a wide range of $\theta$ values, especially in the negative direction, without significantly affecting the objective function value.
+Figure 3: The contour map of the GMM objective function on $\theta-\gamma_{0}$ plane.
+Note: We pick one simulation data where $T=100$ and $\sigma=1$. The GMM objective function is flat around the true values of the intercept of the log marginal cost and the conduct parameter.
+A.5 Additional experiments
+Additional results for different $\sigma$ are shown in Tables 4 and 5. As a summary, the main findings in the main text are robust.
+Table 6 shows the results of the estimation with ado hoc method with different standard deviations of the error terms.
+As in Table 2, the run convergence rate becomes 100% and the accuracy of the conduct parameter estimation is improved.
+However, again, the results with ad hoc improvement do not dominate the results without the improvement for all parameters.
+In Tables 7 and 8, we illustrate that N2SLS with Constraints (12) works for the linear model as in matsumura2023resolving.
 This means that incorporating the conditions is innocuous for the linear model.
-Additional ad hoc improvement is possible by using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14).
-Table 5 shows that the estimation of the conduct parameter $\theta$ improves, that is, the bias is 0.014 and RMSE is 0.217.
-However, the results with ad hoc improvement do not dominate the results without the improvement for all parameters.
+Table 9 demonstrates that relying solely on Constraints (13) and (14) within the N2SLS framework leads to severe bias, implying that only Constraints (13) and (14) cannot resolve the problem.
+Table 10 demonstrates that relying solely on Constraints (12) within the N2SLS framework leads to severe bias—especially in small samples—for parameters other than the conduct parameter, as well as poor convergence.
+While the conduct parameter itself appears stable, this is mechanically due to the imposed domain constraint, which prevents extreme estimates by construction.
+These results underscore that Constraints (12) alone are insufficient and highlight the critical importance of incorporating the equilibrium conditions (13) and (14).
+In large samples, however, the problem is substantially alleviated, as the parameter search becomes less problematic.
 Table 4: Performance comparison ($\sigma=0.5$)
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$ 0.119 4.055 0.346 3.180 0.344 1.497 0.379 1.582
-$\alpha_{1}$ 0.019 0.620 0.054 0.486 0.076 0.429 0.099 0.576
-$\alpha_{2}$ 0.000 0.050 0.003 0.039 0.002 0.029 0.003 0.036
-$\alpha_{3}$ 0.009 0.336 0.025 0.268 0.028 0.129 0.031 0.129
-$\gamma_{0}$ 10.104 11.066 9.397 10.288 13.240 13.280 13.314 13.346
-$\gamma_{1}$0.146 0.340 -0.166 0.264 -0.180 0.198 -0.181 0.192
-$\gamma_{2}$0.072 0.229 -0.079 0.165 -0.087 0.108 -0.086 0.099
-$\gamma_{3}$0.069 0.230 -0.082 0.171 -0.088 0.110 -0.087 0.100
-$\theta$5e+04 1e+05 -3e+04 6e+04 -8e+04 9e+04 -8e+04 9e+04
-Runs converged (%) 100.000 100.000 98.100 98.700
-Sample size ($T$) 100 200 1000 1500
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.006 3.556 0.175 3.158 0.059 0.977 -0.017 0.763
-$\alpha_{1}$0.002 0.545 0.026 0.484 0.009 0.150 -0.003 0.117
-$\alpha_{2}$ 0.002 0.044 0.004 0.037 0.000 0.012 0.000 0.010
-$\alpha_{3}$ 0.001 0.309 0.012 0.266 0.005 0.091 0.000 0.073
-$\gamma_{0}$0.616 3.317 -0.221 2.191 0.275 1.076 0.421 0.972
-$\gamma_{1}$ 0.089 0.399 0.038 0.252 -0.015 0.107 -0.027 0.089
-$\gamma_{2}$ 0.044 0.265 0.022 0.170 -0.007 0.074 -0.011 0.061
-$\gamma_{3}$ 0.046 0.273 0.019 0.178 -0.008 0.078 -0.012 0.059
-$\theta$0.064 0.390 -0.079 0.339 -0.119 0.260 -0.145 0.261
-Runs converged (%) 99.900 100.000 100.000 100.000
-Sample size ($T$) 100 200 1000 1500
+(a)N2SLS without Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ 0.119 4.055 0.346 3.180 0.344 1.497 0.379 1.582 $\alpha_{1}$ 0.019 0.620 0.054 0.486 0.076 0.429 0.099 0.576 $\alpha_{2}$ 0.000 0.050 0.003 0.039 0.002 0.029 0.003 0.036 $\alpha_{3}$ 0.009 0.336 0.025 0.268 0.028 0.129 0.031 0.129 $\gamma_{0}$ 10.104 11.066 9.397 10.288 13.240 13.280 13.314 13.346 $\gamma_{1}$ -0.146 0.340 -0.166 0.264 -0.180 0.198 -0.181 0.192 $\gamma_{2}$ -0.072 0.229 -0.079 0.165 -0.087 0.108 -0.086 0.099 $\gamma_{3}$ -0.069 0.230 -0.082 0.171 -0.088 0.110 -0.087 0.100 $\theta$ -5e+04 1e+05 -3e+04 6e+04 -8e+04 9e+04 -8e+04 9e+04 Runs converged (%) 100.000 100.000 98.100 98.700 Sample size ($T$) 100 200 1000 1500
+(b) N2SLS with Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.006 3.556 0.175 3.158 0.059 0.977 -0.017 0.763 $\alpha_{1}$ -0.002 0.545 0.026 0.484 0.009 0.150 -0.003 0.117 $\alpha_{2}$ 0.002 0.044 0.004 0.037 0.000 0.012 0.000 0.010 $\alpha_{3}$ 0.001 0.309 0.012 0.266 0.005 0.091 0.000 0.073 $\gamma_{0}$ -0.616 3.317 -0.221 2.191 0.275 1.076 0.421 0.972 $\gamma_{1}$ 0.089 0.399 0.038 0.252 -0.015 0.107 -0.027 0.089 $\gamma_{2}$ 0.044 0.265 0.022 0.170 -0.007 0.074 -0.011 0.061 $\gamma_{3}$ 0.046 0.273 0.019 0.178 -0.008 0.078 -0.012 0.059 $\theta$ -0.064 0.390 -0.079 0.339 -0.119 0.260 -0.145 0.261 Runs converged (%) 99.900 100.000 100.000 100.000 Sample size ($T$) 100 200 1000 1500
 Table 5: Performance comparison ($\sigma=2.0$)
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$2.358 11.286 -1.457 8.696 0.231 4.747 0.450 4.028
-$\alpha_{1}$0.360 1.770 -0.227 1.313 0.040 0.744 0.095 0.711
-$\alpha_{2}$0.021 0.138 -0.012 0.142 0.002 0.063 0.006 0.065
-$\alpha_{3}$0.175 0.967 -0.138 0.867 0.006 0.434 0.031 0.347
-$\gamma_{0}$ 9.868 21.289 10.490 13.359 13.151 13.453 13.177 13.384
-$\gamma_{1}$0.015 2.542 -0.130 1.042 -0.202 0.400 -0.186 0.339
-$\gamma_{2}$0.057 1.656 -0.083 0.735 -0.090 0.274 -0.099 0.224
-$\gamma_{3}$0.020 2.115 -0.065 0.719 -0.097 0.279 -0.092 0.232
-$\theta$7e+05 5e+06 -3e+05 2e+06 -1e+05 4e+05 -9e+04 1e+05
-Runs converged (%) 99.600 99.800 98.300 98.900
-Sample size ($T$) 100 200 1000 1500
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$2.008 11.292 -1.149 8.680 0.145 4.655 0.274 3.838
-$\alpha_{1}$0.308 1.773 -0.181 1.313 0.020 0.713 0.041 0.590
-$\alpha_{2}$0.016 0.136 -0.007 0.139 0.003 0.057 0.004 0.043
-$\alpha_{3}$0.147 0.985 -0.113 0.875 -0.003 0.429 0.017 0.333
-$\gamma_{0}$ 0.661 11.910 -1.090 13.475 -0.665 4.501 -0.404 3.245
-$\gamma_{1}$0.044 1.600 0.172 1.770 0.098 0.556 0.066 0.391
-$\gamma_{2}$0.056 1.086 0.088 1.172 0.054 0.346 0.024 0.252
-$\gamma_{3}$0.020 1.205 0.108 1.225 0.049 0.365 0.031 0.267
-$\theta$0.229 0.475 -0.163 0.464 -0.090 0.410 -0.108 0.394
-Runs converged (%) 99.200 99.200 99.700 100.000
-Sample size ($T$) 100 200 1000 1500
+(a) N2SLS without Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -2.358 11.286 -1.457 8.696 0.231 4.747 0.450 4.028 $\alpha_{1}$ -0.360 1.770 -0.227 1.313 0.040 0.744 0.095 0.711 $\alpha_{2}$ -0.021 0.138 -0.012 0.142 0.002 0.063 0.006 0.065 $\alpha_{3}$ -0.175 0.967 -0.138 0.867 0.006 0.434 0.031 0.347 $\gamma_{0}$ 9.868 21.289 10.490 13.359 13.151 13.453 13.177 13.384 $\gamma_{1}$ -0.015 2.542 -0.130 1.042 -0.202 0.400 -0.186 0.339 $\gamma_{2}$ -0.057 1.656 -0.083 0.735 -0.090 0.274 -0.099 0.224 $\gamma_{3}$ -0.020 2.115 -0.065 0.719 -0.097 0.279 -0.092 0.232 $\theta$ -7e+05 5e+06 -3e+05 2e+06 -1e+05 4e+05 -9e+04 1e+05 Runs converged (%) 99.600 99.800 98.300 98.900 Sample size ($T$) 100 200 1000 1500
+(b) N2SLS with Constraints (12), (13), and (14)
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -2.008 11.292 -1.149 8.680 0.145 4.655 0.274 3.838 $\alpha_{1}$ -0.308 1.773 -0.181 1.313 0.020 0.713 0.041 0.590 $\alpha_{2}$ -0.016 0.136 -0.007 0.139 0.003 0.057 0.004 0.043 $\alpha_{3}$ -0.147 0.985 -0.113 0.875 -0.003 0.429 0.017 0.333 $\gamma_{0}$ 0.661 11.910 -1.090 13.475 -0.665 4.501 -0.404 3.245 $\gamma_{1}$ -0.044 1.600 0.172 1.770 0.098 0.556 0.066 0.391 $\gamma_{2}$ -0.056 1.086 0.088 1.172 0.054 0.346 0.024 0.252 $\gamma_{3}$ -0.020 1.205 0.108 1.225 0.049 0.365 0.031 0.267 $\theta$ -0.229 0.475 -0.163 0.464 -0.090 0.410 -0.108 0.394 Runs converged (%) 99.200 99.200 99.700 100.000 Sample size ($T$) 100 200 1000 1500
 Table 6: Ad hoc method using (3) to compute $\varepsilon_{t}^{c}$ and (4) as constraints with Constraints (12), (13), and (14)
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.347 2.978 -0.153 2.547 0.065 0.976 -0.009 0.760
-$\alpha_{1}$0.043 0.452 -0.016 0.388 0.010 0.150 -0.001 0.117
-$\alpha_{2}$0.024 0.065 -0.017 0.054 0.000 0.012 0.000 0.010
-$\alpha_{3}$0.028 0.265 -0.013 0.224 0.005 0.091 0.001 0.073
-$\gamma_{0}$2.458 5.660 -1.258 2.959 -0.132 0.942 -0.153 0.748
-$\gamma_{1}$ 0.112 0.602 0.057 0.260 0.014 0.103 0.014 0.082
-$\gamma_{2}$ 0.053 0.314 0.030 0.172 0.007 0.073 0.010 0.060
-$\gamma_{3}$ 0.061 0.426 0.028 0.181 0.006 0.077 0.008 0.058
-$\theta$ 0.169 0.407 0.109 0.356 -0.003 0.155 0.017 0.116
-Runs converged (%) 100.000 100.000 100.000 100.000
-Sample size ($T$) 100 200 1000 1500
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
-$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
-$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
-$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
-$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
-$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
-$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
-$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
-$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
-Runs converged (%) 100.000 100.000 100.000 100.000
-Sample size ($T$) 100 200 1000 1500
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.835 9.908 -0.561 6.310 0.147 4.424 0.301 3.686
-$\alpha_{1}$0.109 1.563 -0.076 0.935 0.024 0.677 0.047 0.566
-$\alpha_{2}$0.043 0.124 -0.033 0.135 -0.005 0.061 0.001 0.045
-$\alpha_{3}$0.047 0.909 -0.078 0.664 -0.004 0.404 0.018 0.323
-$\gamma_{0}$6.629 20.418 -5.373 16.692 -1.445 4.738 -1.123 3.411
-$\gamma_{1}$ 0.422 2.630 0.384 2.103 0.137 0.580 0.115 0.409
-$\gamma_{2}$ 0.184 1.788 0.185 1.318 0.073 0.357 0.048 0.261
-$\gamma_{3}$ 0.199 1.885 0.215 1.496 0.068 0.376 0.055 0.276
-$\theta$ 0.225 0.459 0.227 0.447 0.098 0.357 0.084 0.313
-Runs converged (%) 99.200 99.800 99.800 100.000
-Sample size ($T$) 100 200 1000 1500
+(a) $\sigma=0.5$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.347 2.978 -0.153 2.547 0.065 0.976 -0.009 0.760 $\alpha_{1}$ -0.043 0.452 -0.016 0.388 0.010 0.150 -0.001 0.117 $\alpha_{2}$ -0.024 0.065 -0.017 0.054 0.000 0.012 0.000 0.010 $\alpha_{3}$ -0.028 0.265 -0.013 0.224 0.005 0.091 0.001 0.073 $\gamma_{0}$ -2.458 5.660 -1.258 2.959 -0.132 0.942 -0.153 0.748 $\gamma_{1}$ 0.112 0.602 0.057 0.260 0.014 0.103 0.014 0.082 $\gamma_{2}$ 0.053 0.314 0.030 0.172 0.007 0.073 0.010 0.060 $\gamma_{3}$ 0.061 0.426 0.028 0.181 0.006 0.077 0.008 0.058 $\theta$ 0.169 0.407 0.109 0.356 -0.003 0.155 0.017 0.116 Runs converged (%) 100.000 100.000 100.000 100.000 Sample size ($T$) 100 200 1000 1500
+(c) $\sigma=2.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.835 9.908 -0.561 6.310 0.147 4.424 0.301 3.686 $\alpha_{1}$ -0.109 1.563 -0.076 0.935 0.024 0.677 0.047 0.566 $\alpha_{2}$ -0.043 0.124 -0.033 0.135 -0.005 0.061 0.001 0.045 $\alpha_{3}$ -0.047 0.909 -0.078 0.664 -0.004 0.404 0.018 0.323 $\gamma_{0}$ -6.629 20.418 -5.373 16.692 -1.445 4.738 -1.123 3.411 $\gamma_{1}$ 0.422 2.630 0.384 2.103 0.137 0.580 0.115 0.409 $\gamma_{2}$ 0.184 1.788 0.185 1.318 0.073 0.357 0.048 0.261 $\gamma_{3}$ 0.199 1.885 0.215 1.496 0.068 0.376 0.055 0.276 $\theta$ 0.225 0.459 0.227 0.447 0.098 0.357 0.084 0.313 Runs converged (%) 99.200 99.800 99.800 100.000 Sample size ($T$) 100 200 1000 1500
 Table 7: N2SLS without Constraints (12) results of the linear model
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.018 0.465 0.007 0.323 -0.008 0.213 -0.006 0.097
-$\alpha_{1}$0.045 2.257 0.024 1.523 0.018 1.016 -0.031 0.455
-$\alpha_{2}$0.001 0.255 -0.001 0.176 -0.004 0.115 0.001 0.051
-$\alpha_{3}$0.005 0.108 0.003 0.075 -0.001 0.050 -0.001 0.022
-$\gamma_{0}$0.061 0.732 -0.005 0.474 -0.021 0.346 -0.005 0.152
-$\gamma_{1}$0.311 3.450 -0.124 1.928 -0.081 1.303 -0.003 0.548
-$\gamma_{2}$ 0.009 0.109 -0.001 0.071 0.003 0.051 0.000 0.023
-$\gamma_{3}$ 0.001 0.108 0.003 0.075 0.003 0.053 0.000 0.022
-$\theta$ 0.047 0.354 0.017 0.209 0.014 0.135 0.003 0.058
-Sample size ($T$) 50 100 200 1000
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.027 1.023 -0.002 0.641 -0.004 0.448 -0.016 0.188
-$\alpha_{1}$0.024 4.396 -0.169 2.965 0.061 2.060 0.011 0.905
-$\alpha_{2}$0.006 0.494 0.016 0.325 -0.007 0.234 -0.006 0.100
-$\alpha_{3}$0.006 0.223 0.002 0.153 -0.001 0.099 -0.003 0.045
-$\gamma_{0}$0.318 1.769 -0.091 1.059 -0.086 0.714 -0.004 0.308
-$\gamma_{1}$ 5.859 210.853 -0.679 6.280 -0.338 2.972 -0.050 1.110
-$\gamma_{2}$ 0.035 0.247 0.011 0.157 0.011 0.104 0.000 0.045
-$\gamma_{3}$ 0.045 0.250 0.012 0.150 0.010 0.104 0.002 0.045
-$\theta$0.399 18.450 0.098 0.738 0.054 0.308 0.009 0.113
-Sample size ($T$) 50 100 200 1000
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.263 2.596 0.071 1.670 -0.040 0.947 -0.002 0.412
-$\alpha_{1}$0.271 10.820 0.008 6.492 0.236 4.263 0.021 1.809
-$\alpha_{2}$0.044 1.253 0.023 0.779 -0.031 0.483 -0.003 0.210
-$\alpha_{3}$0.024 0.584 0.008 0.343 -0.004 0.225 0.003 0.092
-$\gamma_{0}$2.074 19.624 -0.551 3.043 -0.171 1.516 -0.051 0.633
-$\gamma_{1}$ 58.209 1750.688 -2.416 56.909 -3.617 39.044 -0.103 2.334
-$\gamma_{2}$ 0.242 2.430 0.065 0.409 0.020 0.220 0.006 0.093
-$\gamma_{3}$ 0.230 2.328 0.055 0.404 0.010 0.219 0.008 0.092
-$\theta$6.668 233.851 0.372 6.334 0.418 3.820 0.024 0.245
-Sample size ($T$) 50 100 200 1000
-Note: The data generating process follows Matsumura and Otani (2023b).
+(a) $\sigma=0.5$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.018 0.465 0.007 0.323 -0.008 0.213 -0.006 0.097 $\alpha_{1}$ -0.045 2.257 0.024 1.523 0.018 1.016 -0.031 0.455 $\alpha_{2}$ -0.001 0.255 -0.001 0.176 -0.004 0.115 0.001 0.051 $\alpha_{3}$ -0.005 0.108 0.003 0.075 -0.001 0.050 -0.001 0.022 $\gamma_{0}$ -0.061 0.732 -0.005 0.474 -0.021 0.346 -0.005 0.152 $\gamma_{1}$ -0.311 3.450 -0.124 1.928 -0.081 1.303 -0.003 0.548 $\gamma_{2}$ 0.009 0.109 -0.001 0.071 0.003 0.051 0.000 0.023 $\gamma_{3}$ 0.001 0.108 0.003 0.075 0.003 0.053 0.000 0.022 $\theta$ 0.047 0.354 0.017 0.209 0.014 0.135 0.003 0.058 Sample size ($T$) 50 100 200 1000
+(b) $\sigma=1.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.027 1.023 -0.002 0.641 -0.004 0.448 -0.016 0.188 $\alpha_{1}$ -0.024 4.396 -0.169 2.965 0.061 2.060 0.011 0.905 $\alpha_{2}$ -0.006 0.494 0.016 0.325 -0.007 0.234 -0.006 0.100 $\alpha_{3}$ -0.006 0.223 0.002 0.153 -0.001 0.099 -0.003 0.045 $\gamma_{0}$ -0.318 1.769 -0.091 1.059 -0.086 0.714 -0.004 0.308 $\gamma_{1}$ 5.859 210.853 -0.679 6.280 -0.338 2.972 -0.050 1.110 $\gamma_{2}$ 0.035 0.247 0.011 0.157 0.011 0.104 0.000 0.045 $\gamma_{3}$ 0.045 0.250 0.012 0.150 0.010 0.104 0.002 0.045 $\theta$ -0.399 18.450 0.098 0.738 0.054 0.308 0.009 0.113 Sample size ($T$) 50 100 200 1000
+(c) $\sigma=2.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.263 2.596 0.071 1.670 -0.040 0.947 -0.002 0.412 $\alpha_{1}$ -0.271 10.820 0.008 6.492 0.236 4.263 0.021 1.809 $\alpha_{2}$ -0.044 1.253 0.023 0.779 -0.031 0.483 -0.003 0.210 $\alpha_{3}$ -0.024 0.584 0.008 0.343 -0.004 0.225 0.003 0.092 $\gamma_{0}$ -2.074 19.624 -0.551 3.043 -0.171 1.516 -0.051 0.633 $\gamma_{1}$ 58.209 1750.688 -2.416 56.909 -3.617 39.044 -0.103 2.334 $\gamma_{2}$ 0.242 2.430 0.065 0.409 0.020 0.220 0.006 0.093 $\gamma_{3}$ 0.230 2.328 0.055 0.404 0.010 0.219 0.008 0.092 $\theta$ -6.668 233.851 0.372 6.334 0.418 3.820 0.024 0.245 Sample size ($T$) 50 100 200 1000
+Note: The data generating process follows matsumura2023resolving.
 Table 8: N2SLS with Constraints (12) for the linear model
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.013 0.462 0.008 0.322 -0.008 0.213 -0.006 0.097
-$\alpha_{1}$0.096 2.201 0.015 1.511 0.018 1.016 -0.031 0.455
-$\alpha_{2}$ 0.006 0.247 0.001 0.174 -0.004 0.115 0.001 0.051
-$\alpha_{3}$0.004 0.108 0.003 0.074 -0.001 0.050 -0.001 0.022
-$\gamma_{0}$0.054 0.724 -0.002 0.472 -0.021 0.346 -0.005 0.152
-$\gamma_{1}$0.098 2.620 -0.093 1.847 -0.081 1.303 -0.003 0.548
-$\gamma_{2}$ 0.008 0.108 -0.002 0.070 0.003 0.051 0.000 0.023
-$\gamma_{3}$ 0.001 0.107 0.003 0.075 0.003 0.053 0.000 0.022
-$\theta$ 0.023 0.258 0.014 0.197 0.014 0.135 0.003 0.058
-Sample size ($T$) 50 100 200 1000
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$ 0.012 1.015 0.012 0.636 0.001 0.446 -0.016 0.188
-$\alpha_{1}$0.383 4.216 -0.291 2.872 0.015 2.012 0.011 0.905
-$\alpha_{2}$ 0.042 0.466 0.033 0.310 -0.001 0.227 -0.006 0.100
-$\alpha_{3}$ 0.000 0.222 0.004 0.152 0.000 0.099 -0.003 0.045
-$\gamma_{0}$0.233 1.612 -0.067 1.020 -0.076 0.703 -0.004 0.308
-$\gamma_{1}$ 0.406 4.332 -0.024 3.174 -0.178 2.585 -0.050 1.110
-$\gamma_{2}$ 0.025 0.231 0.008 0.154 0.010 0.103 0.000 0.045
-$\gamma_{3}$ 0.034 0.233 0.010 0.146 0.009 0.103 0.002 0.045
-$\theta$ 0.014 0.380 0.018 0.311 0.035 0.260 0.009 0.113
-Sample size ($T$) 50 100 200 1000
-Bias RMSE Bias RMSE Bias RMSE Bias RMSE
-$\alpha_{0}$0.138 2.592 0.142 1.671 0.004 0.936 0.000 0.410
-$\alpha_{1}$0.986 11.090 -0.665 6.325 -0.174 4.042 0.003 1.788
-$\alpha_{2}$ 0.065 1.255 0.112 0.756 0.023 0.447 0.000 0.207
-$\alpha_{3}$0.006 0.589 0.019 0.345 0.003 0.224 0.003 0.092
-$\gamma_{0}$0.123 3.350 -0.298 2.455 -0.081 1.379 -0.047 0.628
-$\gamma_{1}$ 0.635 7.424 0.141 5.668 -0.081 4.255 -0.040 2.159
-$\gamma_{2}$ 0.016 0.484 0.037 0.349 0.009 0.205 0.005 0.093
-$\gamma_{3}$ 0.009 0.517 0.026 0.336 0.000 0.208 0.007 0.092
-$\theta$0.029 0.446 0.025 0.414 0.034 0.381 0.016 0.220
-Sample size ($T$) 50 100 200 1000
-Note: The data generating process follows Matsumura and Otani (2023b).
+(a) $\sigma=0.5$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.013 0.462 0.008 0.322 -0.008 0.213 -0.006 0.097 $\alpha_{1}$ -0.096 2.201 0.015 1.511 0.018 1.016 -0.031 0.455 $\alpha_{2}$ 0.006 0.247 0.001 0.174 -0.004 0.115 0.001 0.051 $\alpha_{3}$ -0.004 0.108 0.003 0.074 -0.001 0.050 -0.001 0.022 $\gamma_{0}$ -0.054 0.724 -0.002 0.472 -0.021 0.346 -0.005 0.152 $\gamma_{1}$ -0.098 2.620 -0.093 1.847 -0.081 1.303 -0.003 0.548 $\gamma_{2}$ 0.008 0.108 -0.002 0.070 0.003 0.051 0.000 0.023 $\gamma_{3}$ 0.001 0.107 0.003 0.075 0.003 0.053 0.000 0.022 $\theta$ 0.023 0.258 0.014 0.197 0.014 0.135 0.003 0.058 Sample size ($T$) 50 100 200 1000
+(b) $\sigma=1.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ 0.012 1.015 0.012 0.636 0.001 0.446 -0.016 0.188 $\alpha_{1}$ -0.383 4.216 -0.291 2.872 0.015 2.012 0.011 0.905 $\alpha_{2}$ 0.042 0.466 0.033 0.310 -0.001 0.227 -0.006 0.100 $\alpha_{3}$ 0.000 0.222 0.004 0.152 0.000 0.099 -0.003 0.045 $\gamma_{0}$ -0.233 1.612 -0.067 1.020 -0.076 0.703 -0.004 0.308 $\gamma_{1}$ 0.406 4.332 -0.024 3.174 -0.178 2.585 -0.050 1.110 $\gamma_{2}$ 0.025 0.231 0.008 0.154 0.010 0.103 0.000 0.045 $\gamma_{3}$ 0.034 0.233 0.010 0.146 0.009 0.103 0.002 0.045 $\theta$ 0.014 0.380 0.018 0.311 0.035 0.260 0.009 0.113 Sample size ($T$) 50 100 200 1000
+(c) $\sigma=2.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.138 2.592 0.142 1.671 0.004 0.936 0.000 0.410 $\alpha_{1}$ -0.986 11.090 -0.665 6.325 -0.174 4.042 0.003 1.788 $\alpha_{2}$ 0.065 1.255 0.112 0.756 0.023 0.447 0.000 0.207 $\alpha_{3}$ -0.006 0.589 0.019 0.345 0.003 0.224 0.003 0.092 $\gamma_{0}$ -0.123 3.350 -0.298 2.455 -0.081 1.379 -0.047 0.628 $\gamma_{1}$ 0.635 7.424 0.141 5.668 -0.081 4.255 -0.040 2.159 $\gamma_{2}$ 0.016 0.484 0.037 0.349 0.009 0.205 0.005 0.093 $\gamma_{3}$ 0.009 0.517 0.026 0.336 0.000 0.208 0.007 0.092 $\theta$ -0.029 0.446 0.025 0.414 0.034 0.381 0.016 0.220 Sample size ($T$) 50 100 200 1000
+Note: The data generating process follows matsumura2023resolving.
+Table 9: N2SLS with Constraints (13) and (14) for the loglinear model
+(a) $\sigma=0.5$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ 0.066 3.646 0.536 3.302 0.638 2.013 0.544 1.606 $\alpha_{1}$ 0.011 0.559 0.084 0.508 0.122 0.451 0.103 0.383 $\alpha_{2}$ 0.001 0.048 0.004 0.040 0.006 0.037 0.006 0.042 $\alpha_{3}$ 0.006 0.317 0.039 0.280 0.049 0.161 0.044 0.131 $\gamma_{0}$ 8.768 9.771 10.695 11.255 12.797 12.854 12.890 12.928 $\gamma_{1}$ -0.143 0.364 -0.173 0.261 -0.182 0.203 -0.180 0.192 $\gamma_{2}$ -0.069 0.247 -0.081 0.164 -0.089 0.111 -0.086 0.099 $\gamma_{3}$ -0.067 0.239 -0.085 0.170 -0.089 0.111 -0.087 0.100 $\theta$ -19426.126 60643.314 -34723.006 57457.978 -64958.969 153375.492 -61625.065 71909.012 Runs converged (%) 100.000 100.000 96.500 93.400 Sample size ($T$) 100 200 1000 1500
+(b) $\sigma=1.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.036 7.004 0.110 5.112 0.576 3.164 0.582 2.252 $\alpha_{1}$ -0.158 1.057 0.022 0.782 0.116 0.682 0.122 0.638 $\alpha_{2}$ -0.009 0.103 0.000 0.102 0.007 0.050 0.008 0.056 $\alpha_{3}$ -0.098 0.615 0.007 0.481 0.036 0.258 0.043 0.195 $\gamma_{0}$ 8.521 15.430 10.972 11.923 12.729 12.840 12.831 12.893 $\gamma_{1}$ -0.039 1.684 -0.184 0.517 -0.184 0.252 -0.176 0.222 $\gamma_{2}$ -0.016 0.977 -0.100 0.348 -0.090 0.154 -0.079 0.127 $\gamma_{3}$ -0.037 0.781 -0.096 0.303 -0.098 0.158 -0.084 0.133 $\theta$ -159519.119 1344515.437 -106885.344 627588.561 -69229.439 166381.207 -62599.755 74618.949 Runs converged (%) 99.800 99.500 97.400 93.000 Sample size ($T$) 100 200 1000 1500
+(c) $\sigma=2.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -2.128 11.259 -1.140 8.691 0.744 5.721 0.778 4.198 $\alpha_{1}$ -0.325 1.767 -0.175 1.309 0.147 1.100 0.142 0.709 $\alpha_{2}$ -0.019 0.139 -0.007 0.145 0.011 0.076 0.010 0.066 $\alpha_{3}$ -0.161 0.971 -0.112 0.866 0.040 0.470 0.057 0.358 $\gamma_{0}$ 9.831 18.315 10.885 13.580 12.904 13.202 12.878 13.060 $\gamma_{1}$ -0.065 2.134 -0.125 1.066 -0.207 0.404 -0.179 0.330 $\gamma_{2}$ -0.100 1.292 -0.082 0.723 -0.090 0.275 -0.095 0.219 $\gamma_{3}$ -0.050 1.475 -0.051 0.759 -0.099 0.281 -0.088 0.229 $\theta$ -696456.922 5688252.310 -418083.860 2991591.323 -84373.063 140028.575 -80810.876 148172.098 Runs converged (%) 99.100 99.100 95.500 90.600 Sample size ($T$) 100 200 1000 1500
+Table 10: N2SLS with Constraints (12) for the loglinear model
+(a) $\sigma=0.5$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.087 3.688 0.145 3.117 0.027 0.975 -0.066 0.764 $\alpha_{1}$ -0.014 0.565 0.021 0.477 0.004 0.150 -0.010 0.117 $\alpha_{2}$ 0.001 0.045 0.003 0.036 0.000 0.012 -0.001 0.010 $\alpha_{3}$ -0.004 0.317 0.009 0.263 0.002 0.090 -0.004 0.073 $\gamma_{0}$ -0.679 3.764 -0.257 2.192 0.015 0.984 0.040 0.815 $\gamma_{1}$ 0.097 0.465 0.041 0.253 0.004 0.104 0.000 0.084 $\gamma_{2}$ 0.049 0.290 0.023 0.170 0.002 0.073 0.003 0.060 $\gamma_{3}$ 0.050 0.288 0.021 0.178 0.001 0.077 0.001 0.058 $\theta$ -0.063 0.389 -0.068 0.333 -0.043 0.199 -0.032 0.167 Runs converged (%) 99.500 99.700 100.000 100.000 Sample size ($T$) 100 200 1000 1500
+(b) $\sigma=1.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.922 8.603 -0.068 5.116 0.037 2.035 0.000 1.556 $\alpha_{1}$ -0.299 1.314 -0.010 0.785 0.005 0.312 0.000 0.240 $\alpha_{2}$ -0.013 0.104 -0.002 0.063 0.001 0.024 0.000 0.019 $\alpha_{3}$ -0.165 0.774 -0.007 0.472 -0.004 0.185 -0.001 0.152 $\gamma_{0}$ -1.767 14.394 -1.001 6.530 -0.208 1.993 -0.156 1.566 $\gamma_{1}$ 0.255 1.949 0.132 0.838 0.034 0.229 0.027 0.174 $\gamma_{2}$ 0.125 1.097 0.053 0.475 0.017 0.150 0.019 0.119 $\gamma_{3}$ 0.099 0.903 0.062 0.481 0.007 0.149 0.014 0.120 $\theta$ -0.098 0.441 -0.060 0.421 -0.061 0.319 -0.058 0.281 Runs converged (%) 98.100 98.700 100.000 100.000 Sample size ($T$) 100 200 1000 1500
+(c) $\sigma=2.0$
+Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -4.309 16.220 -2.888 12.102 -0.143 5.650 0.085 3.693 $\alpha_{1}$ -0.657 2.489 -0.448 1.840 -0.024 0.874 0.012 0.568 $\alpha_{2}$ -0.035 0.178 -0.023 0.170 0.001 0.062 0.002 0.042 $\alpha_{3}$ -0.280 1.174 -0.250 1.181 -0.027 0.545 0.003 0.323 $\gamma_{0}$ 2.206 15.660 -0.576 13.314 -0.897 4.882 -0.646 3.317 $\gamma_{1}$ -0.255 2.096 0.100 1.753 0.119 0.607 0.086 0.399 $\gamma_{2}$ -0.163 1.316 0.045 1.113 0.065 0.363 0.034 0.257 $\gamma_{3}$ -0.155 1.669 0.068 1.127 0.061 0.386 0.041 0.271 $\theta$ -0.201 0.474 -0.142 0.463 -0.053 0.400 -0.050 0.374 Runs converged (%) 98.100 97.200 99.200 99.400 Sample size ($T$) 100 200 1000 1500